### v1

# Searching for local associations while controlling the false discovery rate

Abstract
We introduce local conditional hypotheses that express how the relation between explanatory variables and outcomes changes across different contexts, described by covariates. By expanding upon the model-X knockoff filter, we show how to adaptively discover these local associations, all while controlling the false discovery rate. Our enhanced inferences can help explain sample heterogeneity and uncover interactions, making better use of the capabilities offered by modern machine learning models. Specifically, our method is able to leverage any model for the identification of data-driven hypotheses pertaining to different contexts. Then, it rigorously test these hypotheses without succumbing to selection bias. Importantly, our approach is efficient and does not require sample splitting. We demonstrate the effectiveness of our method through numerical experiments and by studying the genetic architecture of Waist-Hip-Ratio across different sexes in the UKBiobank.
Keywords: Conditional independence; false discovery rate; heterogeneous effects; interactions; knockoffs; selective inference; subgroup analysis.
1 Introduction
1.1 Motivation
As data sets become more detailed and extensive, the methods and models used for analysis must similarly advance in sophistication. A compelling example comes from precision medicine, where the goal is to tailor disease prevention and treatment to individual differences in genes, environments, and lifestyles, departing from a one-size-fits-all approach.
Anna Karenina’s opening line—“All happy families are alike; each unhappy family is unhappy in its own way”—is often cited to describe the heterogeneity of disease. Increasingly, we understand that similar symptoms leading to the same diagnosis can arise from distinct disruptions in biological pathways. Identifying the underlying causes of these symptoms may be the key to achieving the most effective and safest treatments. Tolstoy’s quote is especially fitting, as it refers not to an “unhappy individual” but to an “unhappy family”—suggesting that making broad generalizations requires finding a homogeneous group of subjects. This highlights one of the greatest challenges in handling heterogeneous data: identifying subgroups within which a consistent mechanism operates. This paper presents a specific formulation of this problem and a corresponding solution.
Specifically, we focus on settings where we have an outcome $Y$, a large number $p$ of possible explanatory (or “treatment”) variables $X$, and a set of covariates $Z$ that capture essential individual features which may help identify relatively homogeneous groups within an otherwise heterogeneous population. In other words, we say that these covariates $Z$ define distinct “environments” across which the observations may differ significantly. The goal is to identify the subset of $X$ variables that carry unique information on $Y$, while accounting for the possibility that these key variables may vary across environments.
This situation is quite common in genomics. For instance, in genome-wide association studies (GWAS), the goal is to identify which of the $p$ single nucleotide polymorphisms (SNPs) in $X$ predict a trait $Y$, while controlling for covariates $Z$ such as sex, age, medication use, diet, or healthcare access. Here, different combinations of $Z$ values create diverse environments that may reveal unique genetic associations relevant to each subgroup.
1.2 Main Contributions and Outline
The first contribution of this paper is to introduce a highly expressive family of local conditional independence hypotheses, designed to rigorously address, within a versatile non-parametric statistical framework, the challenge of identifying meaningful associations that may be specific to certain subgroups of a heterogeneous population. This extends the classical notion of population-wide conditional independence, as studied for example by Candès et al. (2018), which can be viewed as a more restrictive “one-size-fits-all” version of our approach. As we will demonstrate, the local hypotheses we propose to study allow for more nuanced and informative insights in the analysis of heterogeneous data and, in certain contexts, provide an intuitive, non-parametric definition of “statistical interaction”.
The second contribution is an extension of the model-X knockoff filter (Candès et al., 2018), originally developed for testing population-wide hypotheses, to enable testing of our local hypotheses while controlling the false discovery rate (FDR). We call our method the Local Knockoff Filter (LKF). A key innovation of the LKF is its ability to test not only pre-defined, fixed hypotheses but also data-driven local hypotheses, automatically identified by powerful machine learning models as likely to be informative for the data at hand.
This approach introduces a new way to efficiently partition the information in a dataset, allowing both learning and testing (Tian and Taylor, 2018; Rasines and Young, 2023; Leiner et al., 2023). Thus, our method can be seen as a “data-thinning” approach (Dharamshi et al., 2024), specifically designed for knockoff inference and taking advantage of the unique capacity of knockoffs to partially “mask” data (Spector and Fithian, 2022).
After outlining the general framework of the LKF, which is very flexible, we describe specific implementations, including an approach tailored to the analysis of large-scale GWAS data. The effectiveness of our approach is then demonstrated through numerical experiments with simulated data, and with an analysis of the genetic basis of waist-hip ratio and its variation between women and men in the UK Biobank data (Bycroft et al., 2018).
1.3 Related Work
The analysis of heterogeneous data extends beyond the focus of this paper and has attracted interest for various reasons. While a comprehensive review is beyond the scope of this work, it is worth noting that heterogeneous data analysis can help mitigate collinearity issues (Peterson et al., 2019), improve the robustness of predictive models to distribution shifts (Berg et al., 2019), and help identify causal associations (Peters et al., 2016).
This paper focuses on identifying subgroups of a heterogeneous population within which there are relatively homogeneous conditional associations. There are many algorithms and models that can discover subgroups of individuals exhibiting coherent behaviors (Novak et al., 2009), detect interactions (Bien et al., 2013; Wen and Stephens, 2014; Tibshirani and Friedman, 2020), or estimate heterogeneous effects in regression and causal inference (Athey and Imbens, 2016; Moore et al., 2019; Künzel et al., 2019; Hahn et al., 2020; Cui et al., 2023); see also Caron et al. (2022) for a review. While such methods are usually not designed to obtain statistical inferences with finite-sample guarantees, they provide powerful tools that can be leveraged in the “learning” phase of our method.
This work is also inspired by the literature on explainability in machine learning; e.g., see Christoph (2020) for a review. A central challenge in this field is enhancing the interpretability of opaque predictive models by identifying “important” variables. There are many definitions of variable importance, from descriptive measures capturing a model’s input-output relationship to inferential measures offering insights into the underlying population. Given the nonlinear nature of many machine learning models, assessments of variable importance are often context-dependent. Despite limitations in current approaches (Ghassemi et al., 2021; Mase et al., 2024), this research highlights the value of “local” importance measures in capturing data heterogeneity.
A few prior studies have addressed heterogeneity in knockoff-based analyses (Barber and Candès, 2015), particularly within the model-X framework (Candès et al., 2018) used in this paper. Gimenez and Zou (2019) introduced a notion of local variable importance expressed as statistical hypotheses testable via the knockoff filter. Their approach formulates inference in terms of conditional associations within neighborhoods in the space of $X$, with each neighborhood corresponding to distinct hypotheses tested under FDR control. To enable efficient inference, these neighborhoods are defined (possibly adaptively) using both original features and synthetic knockoffs, which can complicate interpretability. Additionally, Gimenez and Zou (2019) note that testing all local hypotheses within fixed neighborhoods may significantly reduce power due to smaller effective sample sizes.
In contrast, we use knockoffs only as analytical tools, without influencing the definition or interpretation of the hypotheses. Further, our method adjusts the “locality” of hypotheses in a data-driven manner, striking an effective balance between power and specificity.
Finally, our work is inspired by Li et al. (2022), which extends the knockoff filter to analyze heterogeneous data. While their definition of local conditional hypotheses informs our approach, we address a different problem. Li et al. (2022) assumes each observation belongs to a predetermined environment—for instance, in a GWAS, individuals may be categorized by self-reported race/ethnicity, with these labels representing distinct “environments.”
Our approach differs in two key ways. First, we do not assume fixed environments for individuals. Instead, we allow each variable to “interact” with different covariates, creating subgroup-specific associations. For example, one SNP might interact with sex, defining male and female environments, while another interacts with age, defining younger and older environments. Second, unlike Li et al. (2022), we do not assume the relevant environments for each variable are known in advance.
2 Methods
2.1 Review of Global Conditional Hypotheses
Consider $n$ observations of a vector of random variables $(X,Y,Z).$ For each individual $i\in[n]=\{1,\ldots,n\}$, ${X^{i}=(X^{i}_{1},\ldots,X^{i}_{p})\in\mathbb{R}^{p}}$ describes $p$ explanatory variables, $Y^{i}\in\mathbb{R}$ is the outcome of interest (either numerical or categorical), and $\smash{Z^{i}=(Z^{i}_{1},\ldots,Z^{i}_{m})\in\mathcal{Z}\subseteq\mathbb{R}^{m}}$ represents $m$ covariates. The distinction between explanatory variables and covariates reflects our inferential question: we want to discover which of the $X_{j}$ for $j\in[p]$ is “important” for $Y$, while $Z$ is used to account for known possible confounders.
One way of formalizing this problem, introduced in Candès et al. (2018), is to test under FDR control the following family of global conditional independence hypotheses:
$\mathcal{H}_{j}:Y\perp\!\!\!\perp X_{j}\mid X_{-j},Z\qquad j\in[p],$ (1)
where $X_{-j}$ indicates all the variables in $X$ except $X_{j}$. The term “global” is adopted in this paper to emphasize these hypotheses describe relations between variables across the entire population, as opposed to the more specific local hypotheses that we will introduce later. We refer to Appendix A1 for a brief review of the global knockoff filter method developed by Candès et al. (2018), which is designed to test (1) while controlling the FDR.
Testing (1) can yield discoveries that are both precise and informative. In the GWAS context, it makes it possible to identify a collection of non-redundant SNPs that capture the genetic signature of a trait, minimizing spurious associations (variants that are merely “guilty by association”) and focusing on more likely causal mechanisms. We refer to Appendix A1 for further details on the interpretation of (1) and its connections to parametric inference in generalized linear models and causal inference.
The non-parametric nature of (1) brings two advantages. First, this type of hypothesis can be interpreted without relying on assumptions whose alignment with reality may be difficult to validate, since it emphasizes testing general relationships between variables rather than estimating parameters within a model. Second, this non-parametric approach offers some robustness to heterogeneity, as explained below, even though it is not specifically designed to shed light on the precise nature of heterogeneous relationships.
Consider for example a simple mixture population with $p=2$ and $m=1$, with $Z_{1}=1$ for half of the individuals and $Z_{1}=0$ for the others, indicating two separate environments. Suppose $X_{1}$ and $X_{2}$ are independent standard normal and $Y\sim\mathcal{N}(Z_{1}X_{2}\beta_{2},1)$, for some coefficient $\beta_{2}\in\mathbb{R}$. The hypotheses in (1) are well defined, with $\mathcal{H}_{1}$ true and $\mathcal{H}_{2}$ true if and only if $\beta_{2}=0$. Thus, testing (1) is already useful for distinguishing the importance of $X_{2}$ in predicting $Y$ from the irrelevance of $X_{1}$. However, one limitation of this approach, which we address in this paper, is that it does not inform about how the importance of $X_{2}$ varies with changes in the value of $Z_{1}$. In other words, testing (1) cannot tells us that $X_{2}$ is only locally important among the subgroup with $Z_{1}=1$.
As a first step towards capturing heterogeneity, Li et al. (2022) extended the framework of Candès et al. (2018) to test a family of environment-level hypotheses defined as
$\mathcal{H}^{(e)}_{j}:Y^{(e)}\perp\!\!\!\perp X^{(e)}_{j}\mid X^{(e)}_{-j},Z^{(e)},$ (2)
simultaneously for different environments $e\in\{1,2,\ldots,E\}$. Here, each environment corresponds to a distinct distribution over $(X,Y,Z)$, which may be interpreted as representing a fixed and pre-defined subgroup. This approach, for example, allows testing the importance for a trait of each SNP across distinct ancestries, like African or European.
However, Li et al. (2022) do not address our problem: their approach lacks flexibility to test hypotheses corresponding to different subgroups for different variables and to enable the discovery of relevant subgroups adaptively. We thus present a novel approach to overcome these challenges, beginning with the introduction of a convenient non-parametric framework for capturing data heterogeneity and a new family of local conditional hypotheses.
2.2 A Model-X Framework for Heterogeneous Data
Let us denote as $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{Y}\in\mathbb{R}^{n\times 1}$, and $\mathbf{Z}\in\mathbb{R}^{n\times m}$ the data matrices collecting the observations for all $n$ individuals. To capture heterogeneity without an explicit parametric model, we assume that the joint distribution of the data can be written as:
$P(\mathbf{X},\mathbf{Y},\mathbf{Z})$ $=\prod_{i=1}^{n}p_{X,Z}(X^{i},Z^{i})\cdot p^{i}_{Y\mid X,Z}(Y^{i}\mid X^{i},Z^{i}),$ (3)
where each $p^{i}_{Y\mid X,Z}$ denotes the (unknown) conditional distribution of the outcome $Y^{i}$ for the $i$-th individual, given that individual’s variables and covariates. This distribution may vary across individuals, who are assumed to be independent. To simplify notation, we assume that $p_{X,Z}$ remains constant, though this assumption could be relaxed. The concept of an individual-specific distribution $p^{i}_{Y\mid X,Z}$ is abstract and provides limited opportunities for learning, as each individual essentially forms their own category. We use this framework to capture unknown heterogeneity, but in practice, learning will be conducted at the group level by aggregating information from “similar” individuals, as it will soon become clear.
Given the distribution (3) one may consider testing a family of individual-level conditional hypotheses defined as:
$\mathcal{H}^{i}_{j}:Y^{i}\perp\!\!\!\perp X^{i}_{j}\mid X^{i}_{-j},Z^{i},\qquad i\in[n],j\in[p].$ (4)
Intuitively, this asks whether the distribution $p^{i}_{Y\mid X,Z}$ for the $i$-th individual in (3) depends on $X_{j}$, representing an extreme departure from the global hypotheses of Candès et al. (2018). While these hypotheses are theoretically appealing because they allow us to think of conditional associations on a per-person basis, they are impractical to test within any non-parametric framework, as each hypothesis applies only to a single data point.
2.3 Local Conditional Hypotheses
To gather strength from multiple data points, we introduce a new family of local conditional hypotheses, offering a pragmatic middle ground between the highly flexible but impractical individual hypotheses in (4) and the more rigid environment-level hypotheses in (2). Intuitively, the idea is to define flexible environments specific to each variable $j\in[p]$, grouping together individuals who are relatively homogeneous in the conditional association between $Y$ and $X_{j}$. By testing conditional hypotheses within these local environments, we can pool information from similar individuals, gaining power to make discoveries.
To make this idea precise, we use a collection of partition functions $\nu_{j}:\mathcal{Z}\mapsto[L_{j}]$ that, for each $j\in[p]$, use the covariate space $\mathcal{Z}$ to identify $L_{j}$ disjoint neighborhoods $\ell\in[L_{j}]$, each defining a sub-population of relevance for variable $X_{j}$. For example, in the case of the toy model from Section 2.1, the most informative partition function is $\nu_{1}(Z)=1$ (a single trivial neighborhood encompassing the whole sample, since $X_{1}$ is unimportant for all individuals) and $\nu_{2}(Z)=1+\mathbbm{1}\left[Z_{1}=1\right]$, where $\mathbbm{1}\left[\cdot\right]$ is the indicator function and identifies the two subgroups defined by $Z_{1}=0$ and $Z_{1}=1$, across which $X_{2}$ has different impacts.
For a given collection of partition functions $\nu=(\nu_{1},\ldots,\nu_{p})$, we are interested in testing the family of local conditional hypotheses defined as:
$\mathcal{H}_{j,\ell}(\nu)=\bigcap_{i\in[n]\,:\,\nu_{j}(Z^{i})=\ell}\mathcal{H}^{i}_{j},\qquad j\in[p],\ell\in[L_{j}],$ (5)
where $\mathcal{H}^{i}_{j}$ is the individual-level hypothesis defined in (4). In words, $\mathcal{H}_{j,\ell}(\nu)$ states that $Y$ does not depend on $X_{j}$ conditional on $X_{-j}$ for any of the individuals in the group defined by $\nu_{j}(Z)=\ell$. Considering (5) for all $j\in[p]$ and $\ell\in[L_{j}]$ gives a multiple testing problem with a total of $L=\sum_{j=1}^{p}L_{j}$ hypotheses, if the $\nu_{j}$ are fixed. In the following, we will denote the subset of true null local hypotheses as $\mathcal{H}_{0}(\nu)=\{\mathcal{H}_{j,\ell}(\nu):j\in[p],\ell\in[L_{j}],\mathcal{H}_{j,\ell}(\nu)\text{ is true}\}$. To simplify notation, rather than referring to the number of neighborhoods $L_{j}$ that vary across different variables, we will often use $L_{\max}=\max_{j}(L_{j})$ with the understanding that this is an upper bound for the number of neighborhoods.
Let us take a moment to clarify the interpretation the local hypotheses and how they differ from the population-wide conditional hypotheses in (1) and the environment-level hypotheses in (2). If $L=1$, (5) essentially has the same interpretation as (1). In general, however, a rejection of (5) is more informative than one of (1) because it pinpoints a specific subgroup of individuals displaying a significant association. Further, the connection with the approach of Li et al. (2022) can be seen by noting that $\mathcal{H}_{j,\ell}(\nu)$ in (5) can be interpreted as being equivalent to an environment-level hypothesis
¯$\bar{\mathcal{H}}_{j,\ell}(\nu):Y^{(j,\ell)}\perp\!\!\!\perp X^{(j,\ell)}_{j}\mid X^{(j,\ell)}_{-j},Z^{(j,\ell)}\;\;\;\;j\in[p],\ell\in[L_{j}],$ (6)
where $\smash{(X^{(j,\ell)},Y^{(j,\ell)},Z^{(j,\ell)})}$ denotes a random sample from an “environment” where the distribution of $(X,Y,Z)$ is conditioned on $\nu_{j}(Z)=\ell$. Thus, rejecting $\mathcal{H}_{j,\ell}(\nu)$ helps pinpoint the effect of $X_{j}$ to the sub-population with $\nu_{j}(Z)=\ell$. This highlights the higher flexibility of (5) relative to (2): it allows us to focus on different environments for different variables.
In the toy example in Section 2.1, the fact that $\mathcal{H}^{i}_{2}$ is true if and only if $Z^{i}_{1}=0$ reflects that the local importance of $X_{2}$ for $Y$ depends on the value of $Z_{1}$. In this sense, the hypotheses in (5) provide a non-parametric extension of the concept of interaction between $X_{2}$ and $Z_{1}$, moving beyond the common but somewhat narrow definition as a “departure from additivity in a linear model on a selected scale of measurement” (Wang et al., 2011).
This being said, it is important to emphasize two elements that distinguish testing (5) from the more standard search for statistical interactions within a generalized linear model. First, in our non parametric framework, there is no notion of “effect size”: a variable is either conditionally independent of the outcome, under $\smash{P^{(j,\ell)}_{X,Y,Z}}$, or it is not. Consider a modification of our running example that sets $Y\sim\mathcal{N}((Z_{1}+1)X_{2}\beta_{2},1)$. Now $X_{2}$ is important for the entire population, just with a different effect size in the neighborhoods defined by $Z_{1}=0$ and $Z_{1}=1$. In this case, both $\mathcal{H}_{2,1}$ and $\mathcal{H}_{2,2}$ are true. The hypotheses (5) then can only describe what are sometimes called “sufficient cause interactions” (Weele and Knol, 2014): situations where the outcome occurs when both the “interacting” exposures are present. Interestingly, these type of interactions often lend themselves to a robust, “mechanistic” interpretation, less dependent of a specific scale and linear model. Second, it is important to underscore that failing to reject $\mathcal{H}_{1,2}$ is not equivalent to establishing it is true. The intrinsic asymmetry of the hypothesis testing framework always has to be kept in mind when interpreting the results obtained by testing $\mathcal{H}_{2,1}$ and $\mathcal{H}_{2,2}$.
We conclude this section by highlighting that, crucially, the hypothesis in (5) is truly more informative than (1) only if the partition function $\nu_{j}$ identifies neighborhoods which are relatively homogeneous in the associations between $X_{j}$ and $Y$. Since this information is typically not available a priori, the selection of $\nu$ should generally be data-driven. As mentioned in Section 1, there exist many algorithms that could be applied to learn a meaningful partition function $\nu$ from the data. However, one must be careful that allowing the hypotheses to be random may create the risk of selection bias in the subsequent tests. In particular, we will see that if the same data utilized to select the subgroups were re-used naively to test the hypotheses in (5), the type-I errors might be inflated (Wager and Athey, 2018). Sample splitting could circumvent this issue, but it is wasteful. This paper presents an efficient solution that overcomes the limitations of both aforementioned benchmarks.
To present our method, we begin in the next section by outlining our key modeling assumptions, including the availability of suitable knockoff variables. In Section 2.5, we introduce a knockoff filter to test local hypotheses $\mathcal{H}_{j,\ell}(\nu)$ defined by a fixed partition function $\nu$ while controlling the FDR. Then, in Section 2.6, we extend this method to accommodate a data-driven partition function $\hat{\nu}$, ensuring FDR control for the resulting random hypotheses $\mathcal{H}_{j,\ell}(\hat{\nu})$ conditional on the choice of $\hat{\nu}$.
2.4 Modeling Assumptions and Knockoffs
In general, Equation (3) can be factored as $P(\mathbf{X},\mathbf{Y},\mathbf{Z})=\prod_{i=1}^{n}p_{Z}(Z^{i})\cdot p_{X\mid Z}(Z^{i}\mid X^{i})\cdot p^{i}_{Y\mid X,Z}(Y^{i}\mid X^{i},Z^{i})$. Following a model-X approach (Candès et al., 2018), we treat $\mathbf{Z}$ as fixed and $P_{X\mid Z}$ as known, while allowing $p^{i}_{Y\mid X,Z}$ to be arbitrary and unknown for all $i\in[n]$.
The assumption that $P_{X\mid Z}$ is known may not be justified in every application, but its validity is testable (Romano et al., 2020) and is well-suited for the analysis of data from randomized experiments (Nair and Janson, 2023) or GWAS (Sesia et al., 2019).
The key ingredient of our method are the knockoff variables, which we denote as $\tilde{X}$. As prescribed by Candès et al. (2018), these are synthetic variables constructed by the statistician, using knowledge of $P_{X\mid Z}$, as a function of $X$ and $Z$ without looking at $Y$, so that $\tilde{X}\perp\!\!\!\perp Y\mid X,Z$. Each observation of $X$ is assigned a corresponding knockoff $\tilde{X}$, with $\tilde{\mathbf{X}}\in\mathbb{R}^{n\times p}$ denoting the corresponding data matrix of knockoff variables for $\mathbf{X}$.
Knockoffs are constructed to be exchangeable with $X$ conditional on $Z$. That is, if $[\mathbf{X},\tilde{\mathbf{X}}]\in\mathbb{R}^{n\times 2p}$ is the matrix obtained by concatenating $\mathbf{X}$ with $\tilde{\mathbf{X}}\in\mathbb{R}^{n\times p}$ and, for any $j\in[p]$, $[\mathbf{X},\smash{\tilde{\mathbf{X}}]_{\mathrm{swap}(j)}}$ is obtained by swapping the $j$-th columns of $\mathbf{X}$ and ${\tilde{\mathbf{X}}}$, then
[…]
Constructing knockoffs that satisfy (7) not only requires knowledge of $P_{X\mid Z}$ but it can also be computationally expensive in general. However, the problem is well-studied (Candès et al., 2018; Gimenez et al., 2019; Bates et al., 2020), with practical and effective solutions available for GWAS data (Sesia et al., 2019, 2020, 2021). Therefore, in this paper we assume the availability of suitable knockoffs for the data at hand, using this as a starting point for developing our methodology to test local conditional independence hypotheses.
2.5 A Knockoff Filter for Fixed Local Hypotheses
We now present an extension of the knockoff filter (Candès et al., 2018) that utilizes the augmented data $\mathcal{D}=(\smash{[\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z}})$ to simultaneously test all local conditional independence hypotheses in (5), for a fixed partition function $\nu$, while controlling the FDR.
The first step of our method is to compute a vector of importance scores for all $p$ variables and knockoffs, with separate scores for each region of the covariate space determined by $\nu=(\nu_{1},\ldots,\nu_{p})$. We will denote this vector of scores as $\smash{[\mathbf{T},\tilde{\mathbf{T}}]\in\mathbb{R}^{2{L}}}$, where $L=\sum_{j=1}^{p}{L}_{j}(\nu)$ and ${L}_{j}(\nu)=\max_{z\in\mathcal{Z}}\nu_{j}(z)\in\mathbb{N}$ is the number of disjoint subgroups induced by $\nu_{j}$, for any $j\in[p]$. More precisely, $\mathbf{T}$ (respectively, $\tilde{\mathbf{T}}$) is the concatenation of $p$ sub-vectors $(T_{j,\ell})_{\ell\in[L_{j}]}$ (respectively, $(\tilde{T}_{j,\ell})_{\ell\in[L_{j}]}$) for $j\in[p]$, where each element indexed by $(j,\ell)$ quantifies the importance of $X_{j}$ (respectively, $\tilde{X}_{j}$) in predicting $Y$ within the subgroup defined by $\nu_{j}(z)=\ell$, according to a suitable predictive model for $Y$ given $(X,\tilde{X},Z)$.
To enable provable FDR control, these scores must be computed in such a way that swapping a variable $X_{j}$ with its knockoff $\tilde{X}_{j}$ within any subgroup $\ell\in[L_{j}]$ would have the sole effect of swapping the corresponding scores and in that subgroup, namely $T_{j,\ell}$ and $\tilde{T}_{j,\ell}$. We now formalize this principle within a general framework, which not only encompasses all implementations of our method applied in this paper but also accommodates a range of potential future extensions. This framework is based on a data augmentation idea.
Let $\mathbf{V}\in\{0,1\}^{n\times p}$ be i.i.d. Bernoulli random variables, representing a random noise matrix generated by the analyst independent of everything else. Then, let $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}\in\mathbb{R}^{n\times 2p}$ denote the concatenation of $\mathbf{X}$ and $\tilde{\mathbf{X}}$, with the $i$-th observation of $X_{j}$ swapped with its knockoff if and only if $V_{ij}=1$. Combining $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}$ with $\mathbf{Y}$ and $\mathbf{Z}$ gives us a cloaked data set $\tilde{\mathcal{D}}(\mathbf{V})=(\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z}})$, where the identities of the original variables and knockoffs are hidden. Although $\tilde{\mathcal{D}}(\mathbf{V})$ (deliberately) contains additional noise compared to $\mathcal{D}$, it still carries valuable information about the relation between $X$ and $Y$ because each column of the matrix $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$ contains approximately $n/2$ real observations.
As will soon become evident, the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ is useful for augmenting the data in $\mathcal{D}$, as it can be safely reused across multiple stages of our analysis. In contrast, the remaining information, specifically the true identities of the variables and knockoffs not captured by $\tilde{\mathcal{D}}(\mathbf{V})$, must be used with great care because it can only be accessed once—to compute the test statistics for the corresponding local hypotheses.
Without loss of generality, the importance scores ${[\mathbf{T},\tilde{\mathbf{T}}]}$ can be written as the output of a (possibly randomized) function $\bm{\tau}$ taking as input both $\mathcal{D}$ and $\tilde{\mathcal{D}}(\mathbf{V})$:
$[\mathbf{T},\tilde{\mathbf{T}}]=\bm{\tau}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)}=\big{[}\bm{t}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)},\tilde{\bm{t}}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)}\big{]},$ (8)
where $\bm{t}$ (resp. $\tilde{\bm{t}}$) are defined in terms of $\bm{\tau}$, as its first (resp. last) ${\hat{L}}$ elements. This notation is useful to highlight the distinct roles played by the original dataset $\mathcal{D}$ and the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ in computing importance scores. Specifically, $\tilde{\mathcal{D}}(\mathbf{V})$ can be used freely, while FDR control requires strict constraints on how $\bm{\tau}$ uses the data in $\mathcal{D}$, as explained next.
For any $\mathcal{S}\subseteq[\smash{L}]$, whose elements uniquely identify pairs $(j,\ell)$ for $j\in[p]$ and $\smash{\ell\in[{L}_{j}]}$, let $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathcal{S})}}$ be the matrix obtained from $\smash{[\mathbf{X},\tilde{\mathbf{X}}]}$ after swapping the sub-column $\smash{\mathbf{X}^{(j,\ell)}_{j}}$, which contains all observations of $X_{j}$ in the subgroup $(j,\ell)$, with the corresponding knockoffs, for all $(j,\ell)\in\mathcal{S}$. Similarly, define $\mathcal{D}_{\mathrm{swap}(\mathcal{S})}=([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathcal{S})},\mathbf{Y},\mathbf{Z})$. Then, our importance scores must be computed in such a way that swapping a variable with its knockoff within any subgroup in $\mathcal{D}$ would have the sole effect of swapping the corresponding importance scores for that subgroup. Formally, this means that $\bm{\tau}$ must satisfy:
$\bm{\tau}\big{(}\mathcal{D}_{\mathrm{swap}(\mathcal{S})},\tilde{\mathcal{D}}(\mathbf{V})\big{)}=\big{[}\bm{t}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)},\tilde{\bm{t}}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)}\big{]}_{\mathrm{swap}(\mathcal{S})}.$ (9)
In truth, it is sufficient for (9) to hold in distribution if $\bm{\tau}$ itself is randomized (e.g., it involves cross-validation to tune some hyper-parameters). However, we can safely ignore this detail here to avoid complicating the notation unnecessarily.
One way to achieve (9) is to compute each pair $(T_{j,\ell},\tilde{T}_{j,\ell})$ using an appropriate measure of the importance of $(X_{j},\tilde{X}_{j})$ within a local model $\hat{\psi}^{(j,\ell)}$ predicting $Y$ given $(X,\tilde{X},Z)$. This model can be trained using the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ and a smaller local dataset $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$, including only observations $(X_{j},\tilde{X}_{j},Y,Z)$ for individuals $i\in[n]$ with $\nu_{j}(Z^{i})=\ell$. Figure A2 in Appendix A2 provides a schematic visualization of this workflow. Then, to guarantee (9), it suffices to compute $(T_{j,\ell},\tilde{T}_{j,\ell})$ in such a way that swapping $X_{j}$ and $\tilde{X}_{j}$ in $\mathcal{D}^{(j,\ell)}$ results in $T_{j,\ell}$ being swapped with $\tilde{T}_{j,\ell}$; see Proposition A3 in Appendix A3. With a slight abuse of notation, we write this local symmetry property as:
$T_{j,\ell}\left((\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)},\tilde{\mathcal{D}}(\mathbf{V})\right)$ $=\tilde{T}_{j,\ell}\left((\tilde{\mathbf{X}}_{j},\mathbf{X}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)},\tilde{\mathcal{D}}(\mathbf{V})\right).$ (10)
Notably, (10) is equivalent to the symmetry required, at the whole-sample level, for the importance scores utilized by Candès et al. (2018), and it is very easy to achieve in practice.
Concretely, in this paper, we implement $\hat{\psi}^{(j,\ell)}$ using a regularized generalized linear model (e.g., the lasso), trained with $\mathbf{Y}^{(j,\ell)}$ as the outcome and $[\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V},-j)},\mathbf{Z}]^{(j,\ell)}$ as the predictor matrix, where $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V},-j)}$ denotes the matrix $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}$ without the $j$-th and $(p+j)$-th columns. The additional observations contained in the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ are incorporated into this model through the use of appropriate regularization weights, as explained in Appendix A2. The scores $T_{j,\ell}$ and $\tilde{T}_{j,\ell}$ are then defined as the absolute values of the estimated regression coefficients for the first two (scaled) variables.
While it is convenient to think about the lasso, the main idea behind our approach is applicable with any model. The key to achieving (9) is that the scores $(T_{j,\ell},\tilde{T}_{j,\ell})$ are computed by a local model that sees the other variables (other than $X_{j}$ and $\tilde{X}_{j}$) and the observations in other subgroups (other than $(j,\ell)$) only through the lenses of the swapped data in $\tilde{\mathcal{D}}(\mathbf{V})$. However, revealing the true values of $X_{j}$ and $\tilde{X}_{j}$ within the environment $(j,\ell)$ when fitting the corresponding local model is essential to achieve non-trivial power—otherwise, it would be impossible to tell important variables apart from knockoffs.
Further implementation details on the computation of local importance scores, as well as additional computational shortcuts that enable the efficient application of this method to very large-scale GWAS data, are provided in Appendix A2.
After computing the importance scores $[\mathbf{T},\tilde{\mathbf{T}}]$, we assemble test statistics $\smash{\mathbf{W}\in\mathbb{R}^{{\hat{L}}}}$ for all local hypotheses as in Candès et al. (2018), by computing $\smash{W_{j,\ell}=T_{j,\ell}-\tilde{T}_{j,\ell}}$ for each pair $(j,\ell)$. Finally, we vectorize $\smash{\mathbf{W}}$ and apply the selective SeqStep+ procedure of Barber and Candès (2015) (i.e., the final component of the standard knockoff filter), which computes an adaptive significance threshold determining which hypotheses are rejected. Algorithm 1 summarizes this method, which is guaranteed to control the FDR.
Input: Knockoff-augmented data set $\mathcal{D}=([\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z})$; Random noise matrix $\mathbf{V}$;
Partition function $\nu=(\nu_{1},\ldots,\nu_{p})$; Nominal FDR level $\alpha\in(0,1)$;
1
2Compute the partition width $L=\sum_{j=1}^{p}L_{j}$, where $L_{j}=\max_{z\in\mathcal{Z}}\nu_{j}(z)$;
3 Define local hypotheses $\mathcal{H}_{j,\ell}(\nu)$ as in (5), for $j\in[p]$ and $\ell\in[L_{j}]$;
4 Assemble the cloaked data matrix $\tilde{\mathcal{D}}(\mathbf{V})=([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z})$;
5 for $j\in[p]$ do
6       for $\ell\in[L_{j}]$ do
7             Define the local dataset $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$;
8             Compute importance scores $T_{j,\ell}$ and $\tilde{T}_{j,\ell}$ using the data in $\mathcal{D}^{(j,\ell)}$ and $\tilde{\mathcal{D}}(\mathbf{V})$, in such a way that the local symmetry property in (10) holds; e.g., see Section A2.1 for implementation details;
9             Compute the test statistic $W_{j,\ell}=T_{j,\ell}-\tilde{T}_{j,\ell}$;
10
11Vectorize the test statistic: $\mathbf{W}=(W_{1,1},\ldots,W_{1,L_{1}},W_{2,1},\ldots,W_{p,1},\ldots,W_{p,L_{p}})$;
12 Apply Selective SeqStep+ (Barber and Candès, 2015) at level $\alpha$ to the vector $\mathbf{W}$;
13
Output: List of rejected hypotheses ^$\hat{\mathcal{R}}\subseteq\{\mathcal{H}_{j,\ell}(\nu)\}_{j\in[p],\ell\in[L_{j}]}$.
Algorithm 1 Local Knockoff Filter for fixed hypotheses
Theorem 1.
Algorithm 1 applied to a data set $\mathcal{D}=(\smash{[\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z}})$, where $\tilde{\mathbf{X}}$ are valid knockoffs for $\mathbf{X}$, and a fixed partition function $\nu$, controls the FDR at level $\alpha$ for the local hypotheses $\mathcal{H}_{j,\ell}(\nu)$ defined in (5). That is, $\mathbb{E}[|\hat{\mathcal{R}}\cap\mathcal{H}_{0}(\nu)|/\max(1,|\hat{\mathcal{R}}|)]\leq\alpha$
2.6 A Knockoff Filter for Data-Driven Local Hypotheses
We extend our method to test data-driven local hypotheses (5). The idea is that one can learn an informative partition function $\hat{\nu}=(\hat{\nu}_{1},\ldots,\hat{\nu}_{p})$ using any algorithm and then test the corresponding local hypotheses (5) while controlling the FDR, as long as the learning step only uses the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})=(\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z}})$. In other words, for each $j\in[p]$, we can use any random partition function $\smash{\hat{\nu}_{j}:{\mathcal{Z}}\mapsto{\mathbb{N}}}$ whose parametrization may depend on $\tilde{\mathcal{D}}(\mathbf{V})$; i.e., $\smash{\hat{\nu}_{j}(z;\tilde{\mathcal{D}}(\mathbf{V}))\in[\hat{L}_{j}]}$ for all $z\in\mathcal{Z}$ and some $\hat{L}_{j}\in\mathbb{N}_{+}$.
This two-step framework for adaptively selecting and testing local hypotheses is outlined by Algorithm 2 and visualized schematically in Figure A2 in Appendix A2. We call this approach Adaptive Local Knockoff Filter (aLKF).
Input: Knockoff-augmented data set $\mathcal{D}=([\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z})$; FDR level $\alpha\in(0,1)$;
1 Randomly generate $\mathbf{V}\in\{0,1\}^{n\times p}$, a matrix of i.i.d. Bernoulli noise;
2 Assemble the cloaked data matrix $\tilde{\mathcal{D}}(\mathbf{V})=([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z})$;
3 Looking only at $\tilde{\mathcal{D}}(\mathbf{V})$, learn a data-driven partition function $\hat{\nu}=(\hat{\nu}_{1},\ldots,\hat{\nu}_{p})$, with total width $\hat{L}=\sum_{j=1}^{p}\hat{L}_{j}$; e.g., see Appendix A2.2 for implementation details;
4 Apply Algorithm 1 using the partition function $\hat{\nu}$;
Output: List of rejected hypotheses ^$\hat{\mathcal{R}}\subseteq\{\mathcal{H}_{j,\ell}(\hat{\nu})\}_{j\in[p],\ell\in[\hat{L}_{j}]}$.
Algorithm 2 Adaptive Local Knockoff Filter (aLKF)
aLKF provably controls the FDR for (5) conditional on the selected set of hypotheses, a nontrivial result that does not directly follow from prior work. While the proof strategy draws on Candès et al. (2018) and Li et al. (2022), the data-dependent nature of our hypotheses adds complexity and could introduce selection bias. However, the random noise introduced into the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ protects us from this issue.
Theorem 2.
Algorithm 2 applied to a knockoff-augmented data set $\mathcal{D}=(\smash{[\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z}})$, where $\tilde{\mathbf{X}}$ are valid knockoffs for $\mathbf{X}$, controls the FDR at level $\alpha$ for the family of local hypotheses $\mathcal{H}_{j,\ell}(\hat{\nu})$ in (5) conditional on $\smash{\hat{\nu}}$. That is, $\mathbb{E}[|\hat{\mathcal{R}}\cap\mathcal{H}_{0}(\hat{\nu})|/\max(1,|\hat{\mathcal{R}}|)\mid\hat{\nu}]\leq\alpha$.
A relatively simple yet effective implementation of aLKF uses sparse generalized linear models with interactions (Bien et al., 2013) to learn informative partition functions. While the details of this implementation are deferred to Appendix A2.2, the underlying idea is straightforward, particularly in the case of a single binary covariate $Z_{1}\in\{0,1\}$.
For example, we can fit the lasso to predict $\mathbf{Y}$ using $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$ and $\smash{\mathbf{Z}_{1}}$, including all possible pairwise interactions between $\mathbf{Z}_{1}$ and $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$. For each $j\in[p]$, the partition function $\smash{\hat{\nu}_{j}}$ then divides the samples into two subgroups, corresponding to $\nu_{j}(0)=1$ and $\nu_{j}(1)=1$, if the model selects at least one interaction term for either the $j$-th or $(j+p)$-th variables in $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$. Otherwise, $\smash{\hat{\nu}_{j}(z)}=1$ for all $z$ values. In general, when dealing with $m$ covariates, this approach is applied using a hyper-parameter $G_{\max}\in\mathbb{N}$ controlling the maximum number of interactions that may be selected per variable (e.g., $G_{\max}=2$).
As detailed in Appendix A2.2, this implementation can be extended to handle multiple binary covariates as well as non-binary covariates. Additionally, Appendix A2.3 describes a specialized implementation with computational shortcuts designed for GWAS-like settings, characterized by a very large number $p$ of variables and a modest number $m$ of covariates.
3 Empirical Demonstrations
We demonstrate the performance of aLKF in two simulated settings and one data analysis.
3.1 Simulation: Finding Highly Localized Important Variables
We generate synthetic data with $p=20$ binary variables and $m=80$ covariates, where the first 20 covariates are binary and the remaining 60 are continuous in $[0,1]$. Details on the joint distribution of these variables and the construction of knockoffs are in Appendix A4.
The outcome is generated using a “causal” linear model with heterogeneous effects and homoscedastic standard Gaussian noise. The expected outcome for the $i$-th individual is $\mathbb{E}\left[Y^{i}\mid X^{i},Z^{i}\right]=\sum_{j=1}^{20}X^{i}_{j}\beta_{j}^{i}+\sum_{j=1}^{80}Z^{i}_{j}\gamma_{j}$, where $\smash{\beta_{j}^{i}=\bar{\beta}_{j}Z_{l_{j,1}}^{i}Z_{l_{j,2}}^{i}}$ for some $l_{j,1},l_{j,2}\in[20]$. In words, the effects of $X_{j}$ are subgroup-specific, with $\smash{\beta_{j}^{i}}$ determined by a parameter $\smash{\bar{\beta}_{j}}$ and the interaction of two randomly selected binary covariates $Z_{l_{j,1}},Z_{l_{j,2}}$. For $j\in[10]$, $\smash{\bar{\beta}_{j}=0}$, while $\smash{\bar{\beta}_{j}\in\{\pm 4\}}$ for $j\in\{11,\ldots,20\}$, with independent random signs. Thus, if $X_{j}$ has an effect on $Y$, it has it only in the subgroup where $Z_{l_{j,1}}=1$ and $Z_{l_{j,2}}=1$, comprising approximately 1/4 of individuals when $l_{j,1}\neq l_{j,2}$. The indices $l_{j,1},l_{j,2}$ are sampled with replacement from $[20]$, independently for each $j$. The covariate main effects, captured by $\gamma$, are homogeneous. The first 40 covariates have no effects: $\gamma_{1}=\ldots=\gamma_{40}=0$. Among the remaining 40 covariates, half (randomly chosen) have effects with absolute value 4 and independent random signs, while the others have no effects; i.e., $\gamma_{41},\ldots,\gamma_{80}\in\{0,\pm 4\}$.
Given a sample of $n$ independent observations, our goal is to identify which variables are non-null within which subgroups, aiming for both power (many rejections) and precision (informative subgroups). We apply our method (aLKF) at the nominal FDR level $\alpha=0.1$, using the lasso-based approach for importance scores (Appendix A2.1) and learning data-driven partitions via a lasso model with interactions (Appendix A2.2), allowing at most two interactions per variable ($G_{\max}=2$). We compare aLKF against three benchmarks: (1) Global-KF, the knockoff filter of Candès et al. (2018), which controls the FDR but tests only population-wide hypotheses; (2) Split-LKF, a less powerful version of aLKF that splits the data, using half to learn the partition function and the other half to test local hypotheses, sacrificing power due to reduced sample size but maintaining FDR control; (3) Naive-LKF, an invalid version of aLKF that uses all data twice—to learn the partition and to test local hypotheses—without cloaking, failing to control the FDR due to selection bias.
For each method, we report the false discovery proportion (FDP), the power, and a “homogeneity” metric that captures the informativeness of their findings, averaged over 100 independent experiments. These metrics are defined as follows.
The FDP is defined intuitively from (5): it is the proportion of discoveries for which the variable $X_{j}$ has no causal effect on $Y$ for any individuals within the reported subgroup (which is always equal to the whole sample for the global approach), as determined by the true causal model. Ideally, we aim to keep the average FDP, or FDR, below 10%.
Power is defined as the proportion of variables with a non-zero causal effect that are correctly identified in at least one subgroup, even if the identified subgroup does not precisely match the true subgroup with non-zero effects, provided the rejected variable has an effect in the identified group. Ideally, power should be as close to one as possible.
The homogeneity metric is defined as the proportion of individuals within a reported subgroup for whom a discovered variable $X_{j}$ has a non-zero coefficient in the true causal model. Ideally, homogeneity should be close to one, indicating that the identified subgroups are highly homogeneous in their conditional association between $Y$ and $X_{j}$.
Figure 1 reports the performances of all methods over 100 independent experiments. These results validate our method and demonstrate its advantages relative to the three benchmarks. In summary, the aLKF discoveries are more informative than those of Global-KF, more numerous than those of Split-LKF, and more reliable than those of Naive-LKF.
Figure 1: Performance of the adaptive Local Knockoff Filter (aLKF) and benchmark methods on synthetic data. The informativeness of the discoveries is quantified by the homogeneity of the corresponding subgroups (higher is better). The nominal FDR level is 0.1.
Global-KF is effective at discovering global associations and controls the FDR. However, it cannot identify the subgroups where significant variables have non-zero effects. This limitation is evident in Figure 1, where the homogeneity remains relatively unchanged as the sample size increases. In contrast, aLKF uncovers increasingly informative discoveries as the dataset grows, ultimately enabling precise identification of subgroups with non-zero effects. It avoids the selection bias that inflates the FDR in Naive-LKF and overcomes the inefficiency of Split-LKF, which loses power due to unnecessary sample splitting.
3.2 Simulation: Searching for Gene-Environment Interactions
The following experiments simulate a search for gene-environment interactions in GWAS data, using the real chromosome-21 genotypes of unrelated white non-British subjects in the UK Biobank. We use the knockoffs defined by Sesia et al. (2021). There are $14,733$ individuals in total and $p=8,832$ variants. We randomly sample 30% of the individuals ($n=4,420$) to imitate a more typical high-dimensional GWAS setting.
We simulate two independent binary random covariates $Z_{1}$ and $Z_{2}$, independently across individuals from a Bernoulli($0.5$) distribution. The results of additional experiments using only one covariate can be found in Appendix A6. A continuous outcome $Y$ is simulated based on a linear model: $Y^{i}=\beta_{0}X^{i}+(\beta_{1}X^{i})*\mathbf{1}(Z_{1}^{i}=0)+\epsilon^{i}$, with $\beta_{0},\beta_{1}\in\mathbb{R}^{p}$, where $\epsilon^{i}\sim N(0,1)$ independently for each $i\in[n]$. Thus, $Z_{1}$ is used to create local effects, among the subgroup with $Z_{1}=0$, while $Z_{2}$ is unimportant. We randomly choose 1% of the genetic variants to be non-null, in the sense that they have a non-zero coefficient either globally (in $\beta_{0}$) or locally (in $\beta_{1}$). Additional results obtained using different proportions of non-null variants (0.05% and 2%) can be found in Appendix A6. Each non-null variant is randomly assigned as either a global or a local effect according to the desired proportions.
The nonzero elements in $\beta_{0}$ and $\beta_{1}$ are initialized with absolute values equal to $a/\sqrt{n}$, where the signal amplitude $a\in\mathbb{R}^{+}$ is varied as a control parameter, with independent random signs. Additionally, to mitigate disparities in the detectability of global versus local effects, each element in $\beta_{0}$ and $\beta_{1}$ is re-scaled by a factor equal to the square root of proportion of individuals in the affected subgroup. For global effects, this adjustment factor is 1, while for local effects it is approximately $\sqrt{2}$.
To maintain sufficient power in the presence of strong collinearity among nearby genetic variants on the same chromosome (linkage disequilibrium), we adopt the approach of Sesia et al. (2020, 2021). The genome is partitioned into contiguous blocks of variants in strong linkage disequilibrium, with conditional hypotheses tested at the block level. These blocks serve as the units of inference, replacing individual SNPs. In this simulation, the median width of the blocks is 3 kilobases (kb); see Appendices A2.3.4 and A5 for further details.
Our goal is to identify which variables are non-null within which subgroups, seeking both high power and precision. We apply aLKF as described in Section 3.1, allowing the data-driven partition function to select at most one interaction per variable ($G_{\max}=1$). In addition to the three benchmarks from Section 3.1Global-KF, Split-LKF, and Naive-LKF—we introduce a fourth benchmark, Fixed-LKF, a non-adaptive version of LKF.
Fixed-LKF applies Algorithm 1 using a fixed partition function with four local environments for each variable: $(Z_{1},Z_{2})\in\{0,1\}^{2}$. This does not utilize data cloaking and computes importance scores within each local environment using only the data from that environment, similar to applying Global-KF to approximately one quarter of the observations. While Fixed-LKF tests hypotheses similar to those of aLKF and controls the FDR, it cannot adaptively identify informative partitions. Consequently, it may lose power, due to reduced effective sample sizes, and produce less informative discoveries, making it more difficult to distinguish truly local effects from those that are likely global.
For each method, we report the FDP, power, and “homogeneity” of the discoveries, as in Section 3.1. However, to gain additional insight, we evaluate two separate power metrics: global power and local power. The global power measures the proportion of variables with a global non-zero effect that are correctly identified in at least one subgroup. The local power measures the proportion of variables with a local non-zero effect that are correctly identified in at least one subgroup. As in Section 3.1, these power definitions are agnostic to whether the identified subgroup exactly matches the subgroup with true causal effects.
Figure 2 presents the results of these experiments as a function of the signal amplitude parameter $a$, fixing the proportion of causal effects that are local equal to 50%. The results show that aLKF achieves the highest overall power and the highest power for local effects among all methods with valid FDR control. It has comparable power for global effects to Global-KF but demonstrates significantly higher power for local effects. Both Split-LKF and Fixed-LKF have lower power than aLKF. The informativeness of the discoveries, measured by homogeneity, is highest for aLKF and Fixed-LKF. Note that Fixed-LKF always achieves a homogeneity of 1 by definition for all correct rejections, whereas aLKF can also reject local effects for the entire population. As expected, Naive-LKF fails to control the FDR.
Figure 2: Performance of the adaptive Local Knockoff Filter (aLKF) and benchmark methods on real genotype data with a simulated outcome, as a function of the signal amplitude. Half of the important genetic variables have global causal effects, while the other half have local effects. Other details are as in Figure 1.
Appendix A6 provides additional details, including the heritability associated with the signal amplitudes in Figure 2, and results from experiments conducted under similar setups, with qualitatively similar conclusions. These include experiments varying the proportions of local and global effects or the sparsity levels (0.05% and 2%), and independent analyses performed with Global-KF in each of the four subgroups defined by two binary covariates.
3.3 Data Analysis: Genetic Underpinning of WHR Across Sexes
We study the genetic basis of Waist-Hip Ratio (WHR) across sexes using the UK Biobank data (Bycroft et al., 2018); see Appendix A5 for details. Previous studies have examined sex-specific effects for WHR or WHR adjusted for BMI, identifying between 7 and 44 sex-specific loci (Heid et al., 2010; Randall et al., 2013; Winkler et al., 2015; Shungin et al., 2015; Funkhouser et al., 2020). Given the relatively small number of loci discovered in prior research, we anticipate a high proportion of global effects.
We work with $n=346,119$ unrelated British individuals for whom we have genotypes ($p=591,513$ variants) and measures on sex (46.35% male), age, height, weight, BMI, waist and hip circumference. We define WHR = waist circumference / hip circumference.
We apply aLKF using the specialized implementation described in Appendix A2.3, designed for large-scale GWAS data with a very large number $p$ of explanatory variables. This implementation leverages lasso models and two key computational shortcuts: pre-screening and batch computation of importance scores. The lasso models adjust for the following covariates: age, age squared, the first five genetic principal components, and sex (if applicable). The target FDR level is $\alpha=0.1$. Similar to the simulation in Section 3.2 and as discussed in Appendix A2.3.4 and A5, inference is conducted using blocks of SNPs, with median width equal to 81 kilobases (kb), as the units of inference.
To learn an informative partition function, after pre-screening, we fit a lasso model including all possible pairwise interactions between each of the pre-screened SNPs and sex; see Appendix A2.3.4 for further details on the process. This leads to the definition of 15,001 block-level hypotheses for the full sample and 175 hypotheses each for males and females.
As shown in Table 1, aLKF yields 541 rejections, of which 57 are sex-specific (46 for females and 11 for males). Of the 46 female-specific rejections, 41 are unique to females, while 5 are shared with males. Overall, the number of local rejections for females or males slightly exceeds those reported in prior studies (Heid et al., 2010; Randall et al., 2013; Funkhouser et al., 2020; Winkler et al., 2015), while remaining within a comparable range. The larger number of loci identified for females aligns with previous findings, which observed multiple variants with significant effects in females but smaller or no effects in males (Heid et al., 2010; Randall et al., 2013; Winkler et al., 2015; Funkhouser et al., 2020).
Table 1: Summary of local and global discoveries for WHR on the UK Biobank GWAS data set, obtained with different methods at the 10% FDR level.
Local environment Number of discoveries
aLKF Global-KF Fixed-LKF Separate-KF
All individuals 484 595 NA NA
Female 46 NA 257 271
Male 11 NA 82 63
Total 541 595 339 334
The Manhattan plot in Figure 3 highlights the discoveries made by aLKF, with female-specific findings shown in pink and male-specific findings in blue. Notably, the largest scores are associated with local hypotheses, and the scores for rejected local hypotheses are distributed across the full range, rather than being concentrated at the lower end.
Figure 3: Manhattan plot of aLKF test statistics for the analysis of WHR using the UK Biobank GWAS data. Each dot represents a local hypothesis for either the entire sample or a sex-specific subgroup. The x-axis indicates the genomic location of the leading SNP within each rejected region, labeled by chromosome, while the y-axis shows the corresponding test statistic. The dashed red line marks the rejection threshold for FDR control at 10%. Discoveries specific to females are shown in pink, and those specific to males are in blue.
Table LABEL:tab:gwas-local-annotation in Appendix A7 lists the 57 sex-specific discoveries reported by aLKF, along with validation results. Of these, 28 regions show evidence of sex differences based on prior studies (Heid et al., 2010; Randall et al., 2013; Winkler et al., 2015; Funkhouser et al., 2020). Additionally, 26 regions have been associated with WHR or a closely related phenotype, WHR adjusted for BMI, in prior studies (Sollis et al., 2023; Costanzo et al., 2023), but the discovery of sex differences is novel. Three regions have been weakly associated with WHR or WHR adjusted for BMI in earlier studies (Costanzo et al., 2023).
Table 1 also contrasts the results of aLKF with three benchmark approaches. Global-KF identifies slightly more global discoveries (595 genetic regions), consistent with the experiments in Figure A4 (Appendix A6), which show higher power for Global-KF when most causal effects are global. However, while Global-KF captures 37 of the 57 sex-specific loci identified by aLKF, it misses 20, including 6 previously reported sex-specific effects (see Table LABEL:tab:gwas-local-annotation in Appendix A7). Moreover, Global-KF does not distinguish between sex-specific and global effects, reducing its informativeness in this context.
The Fixed-KF approach described in Section 3.2 results in 339 total rejections. Applying Separate-KF (the global knockoff filter) separately to the male and female datasets (185,694 observations for females, 160,425 for males) yields 271 rejections for females and 63 for males, for a combined total of 334. These findings underscore that aLKF is a powerful method for testing local hypotheses, consistent with earlier numerical experiments.
Finally, the Venn diagram in Figure A13 (Appendix A7) compares the unique loci discovered by aLKF, Global-KF, and Fixed-KF. These approaches agree on 215 loci, while aLKF and Global-KF share 277 additional loci. Table A3 (Appendix A7) further compares the partition-specific rejections between aLKF and Fixed-KF.
4 Discussion
We introduced conditional hypotheses to explore dependencies between explanatory variables and an outcome within contexts defined by other covariates. Our inferential method uses knockoffs to create a “cloaked” version of the data, enabling the identification of heterogeneity while preserving the ability to test local hypotheses after unmasking the knockoffs and original variables. An application to the interaction between sex and DNA variation in the genetic architecture of WHR highlights the potential of this approach.
Applying the LKF necessitates generating knockoffs based on the conditional distribution of $X$ given $Z$, assuming this distribution is known. In our study, this was straightforward since sex is independent of genotypes from non-sexual chromosomes. However, in other contexts, obtaining conditionally valid knockoffs may be more complicated. In such cases, approximate second-order knockoffs may be beneficial (Chu et al., 2024).
The data analysis and simulations presented in this paper all reflect the results of one knockoff generation, with associated randomness. In practical applications, one might want to resort to multiple knockoff to increase the stability of selection (Ren et al., 2023).
The first step of our procedure consists in identifying the relevant “neighborhoods” in which to test for association between $Y$ and each of the $X_{j}$. In the examples considered in this paper, we found that the lasso with interactions can be useful in defining a partition $\nu_{j}(Z)$. We also noted that different penalization levels for main effects and interactions might be appropriate to navigate the trade-off between the precision of the hypotheses we decided to test with the power to reject them, linked to the size of the subgroup defined by $\nu_{j}(Z)$ (see Appendix A2.3.1). A better balance between these might be achieved with a resolution-adaptive procedure as the one described in Gablenz and Sabatti (2024).
The local conditional hypotheses introduced in this paper can be thought as one non parametric definition of “interaction” between explanatory variable $X_{j}$ and covariates $Z$ in influencing outcome $Y$. Extending this idea, one could explore interactions between different explanatory variables $X_{j}$ and $X_{k}$. While some elements of our proposed framework can be adapted for this purpose, careful consideration is needed to properly index and label hypotheses to ensure interpretability. We leave this extension for future work.
Appendix LABEL:app:extension-ralkf, instead, presents an extension of our method designed to detect variables with a robust local association across different subgroups, and demonstrates its practical usefulness in a “transfer learning” setting inspired by Li et al. (2022).
A software implementation of our methods is available at https://github.com/msesia/i-modelx, along with the code needed to reproduce our simulations and data analysis.
Acknowledgments
P. G. and C. S. were supported by NSF grant DMS 2210392. M. S. was supported by NSF grant DMS 2210637. P. G. was supported by a Ric Weiland fellowship. This research has been conducted using data from UK Biobank (Applications 27837 and 74923), a major biomedical database.
Appendix A1 Review of Relevant Technical Background
A1.1 The Interpretation of Global Conditional Hypotheses
The model-X problem studied by Candès et al. (2018) can be thought of as testing, for all $j\in[p]$, whether $X_{j}$ is associated with $Y$ given $Z$ and all variables excluding $X_{j}$ (i.e., $X_{-j}$); that is, whether the null hypothesis defined in (1) holds true: $\mathcal{H}_{0,j}:Y\perp\!\!\!\perp X_{j}\mid X_{-j},Z$. For example, if one assumed a generalized linear model for $Y\mid X,Z$ (which we do not do here), then the hypothesis in (1) would reduce, under relatively mild assumptions on $P_{X\mid Z}$, to stating that the linear coefficient for $X_{j}$ is zero (Candès et al., 2018).
More generally, without any parametric model for the distribution of the outcome, a rejection of (1) can be interpreted as a discovery that $X_{j}$ has some effect on $Y$ conditional on the other observed variables, for at least some of the individuals in the population. (Note that we are not assuming the individuals are identically distributed.) These discoveries can be generally useful to screen variables in high-dimensional data analyses (Candès et al., 2018), to help prioritize follow-up studies (Sesia et al., 2020), and in some cases even to make approximate causal inferences (Li et al., 2022). In fact, under some additional assumptions such as the absence of unmeasured confounders, a rejection of (1) can sometimes be rigorously interpreted as stating that $X_{j}$ causes $Y$ (Bates et al., 2020).
A1.2 The Connection to Causal Inference
To establish an explicit connection between conditional independence testing and causal inference, it would be necessary to introduce two additional assumptions not made in this paper. The first additional assumption would be that the variables $X$ can be interpreted as randomized treatments that may depend on the observed covariates $Z$ but are independent of anything else. The second additional assumption would be that the joint distribution of $(X,Y,Z)$ takes the form of a structural equation model (Bollen, 2014; Pearl, 2009) in which $Z$ and $X$ may cause $Y$, but not the other way around, as visualized by the directed acyclic graph shown in Figure A1.
Figure A1: Graphical representation of a non-parametric causal model linking the treatment ($X$), the outcome ($Y$), the measured covariates ($Z$), and possibly also other unmeasured covariate ($C$). A typical goal in this setting would be to test whether a particular treatment $X_{j}$ has any causal effect on the outcome. The joint distribution of $X\mid Z,C$ is assumed to be known and may depend only on $Z$, so that $X\perp\!\!\!\perp C\mid Z$.
It is easy to see this setup implicitly rules out confounding in a test of conditional independence between $X$ and $Z$, because the treatments are conditionally independent of any other unmeasured covariate $C$. Therefore, to test whether $X_{j}$ has a causal effect on $Y$, it suffices to test whether the conditional independence hypothesis in (1).
Proposition A1 (From Bates et al. (2020)).
Let $C$ be any unmeasured covariate. If $X\perp\!\!\!\perp C\mid Z$, a valid test of the conditional independence null hypothesis $\mathcal{H}_{0,j}$ in (1) is also a valid test of the stronger null hypothesis
$\mathcal{H}_{0,j}^{*}:Y\perp\!\!\!\perp X_{j}\mid Z,X_{-j},C.$ (A1)
Proof.
The proof is analogous to that of Proposition 1 in Bates et al. (2020). Suppose $\mathcal{H}_{0,j}^{*}:Y\perp\!\!\!\perp X_{j}\mid Z,X_{-j},C$ is true. Then, it follows from $C\perp\!\!\!\perp X_{j}\mid Z,X_{-j}$ that $(Y,C)\perp\!\!\!\perp X_{j}\mid Z,X_{-j}$. Therefore, $\mathcal{H}_{0,j}$ must also be true. ∎
A1.3 The Global Knockoff Filter
We briefly review here the global knockoff filter method developed by Candès et al. (2018), which is designed to test the family of global conditional independence hypotheses defined in (1) while controlling the FDR.
Using a knockoff-augmented data set $\mathcal{D}=(\smash{[\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z}})$, with variables $\mathbf{X}$, knockoffs $\tilde{\mathbf{X}}$, outcomes $\mathbf{Y}$, and covariates $\mathbf{Z}$, the global knockoff filter fits a predictive model for $Y$ given $X,\tilde{X}$ and $Z$. This model is utilized compute importance measures $T_{j}$ and $\tilde{T}_{j}$ for each $X_{j}$ and $\tilde{X}_{j}$. Any model can be employed for this purpose, as long as swapping $X_{j}$ with $\tilde{X}_{j}$ only results in $T_{j}$ being swapped with $\tilde{T}_{j}$. Formally, we may write this property as:
$T_{j}\left(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{X}_{-j},\tilde{\mathbf{X}}_{-j},\mathbf{Y},\mathbf{Z}\right)$ $=\tilde{T}_{j}\left(\tilde{\mathbf{X}}_{j},\mathbf{X}_{j},\mathbf{X}_{-j},\tilde{\mathbf{X}}_{-j},\mathbf{Y},\mathbf{Z}\right),$ (A2)
which may be seen as a global counterpart of our local symmetry property defined in (10).
A typical choice is to fit a (generalized) lasso model (Tibshirani, 1996) and define the importance measures as the absolute values of the scaled regression coefficients for $X$ and $\tilde{X}$, after tuning the regularization via cross-validation.
After computing $T_{j}$ and $\tilde{T}_{j}$ for all $j\in p$, these importance scores are combined pairwise into anti-symmetric statistics $W_{j}$; i.e., $W_{j}=T_{j}-\smash{\tilde{T}_{j}}$. This ensures that the signs of the $W_{j}$ are mutually independent coin flips for all $j$ corresponding to a true $\mathcal{H}_{0,j}$, while larger and positive values provide evidence against the null. Letting $\epsilon\in\{-1,+1\}^{p}$ be an independent random vector such that $\epsilon_{j}=+1$ if $\mathcal{H}_{0,j}$ is false and $\mathbb{P}[\epsilon_{j}=+1]=1/2$ otherwise, then it can be proved that $W$ satisfies the flip-sign property:
[…]
where $\odot$ indicates element-wise multiplication. Intuitively, the sign of $W_{j}$ gives rise to a conservative binary p-value $p_{j}$ for $\mathcal{H}_{0,j}$. That is, $p_{j}=1/2$ if $W_{j}>0$ and $p_{j}=1$ otherwise, as long as $\mathcal{H}_{0,j}$ is true.
Finally, a rejection rule for (1) can be obtained by applying a sequential testing procedure to these p-values, in the order defined by the absolute values of $W$. As each of these p-values contains a single bit of information, an appropriate sequential testing procedure is the selective SeqStep+ test of Barber and Candès (2015), which can compute an adaptive threshold for $W_{j}$ controlling the FDR below any desired threshold.
Although inspired by it, the local knockoff filter method developed in this paper is designed to address a problem that differs significantly from the one targeted by the global knockoff filter of Candès et al. (2018). First, our method tests the family of local conditional hypotheses defined in (5), as opposed to the global hypotheses in (1). Second, it treats these hypotheses as random, relying on a data-driven partition function $\hat{\nu}$, rather than treating them as fixed a priori. Technically, the differences in operation lie in how the vector of test statistics $W$ is defined and computed. Once $W$ is obtained, both approaches apply the selective SeqStep+ test of Barber and Candès (2015) in a similar manner.
Appendix A2 Method Implementation
We now describe in more details concrete implementations of our method. We start with an approach to compute powerful importance scores for testing local hypotheses with Algorithm 1. We then introduce a versatile algorithm to learn informative local environments using generalized linear models with interactions. Finally, we describe a specialized implementation of our method that is designed for settings similar to GWAS, where there is a very large number $p$ of possible explanatory variables and a modest number $m$ of covariates.
A2.1 Computing Powerful Local Importance Scores
We describe here in more detail a specific implementation of Algorithm 1 that uses regularized generalized linear models to compute powerful local importance scores satisfying (9). While this is not the only possible approach, it is one that we have found to perform well in practice, particularly when the dataset size is moderate. Additionally, it serves to illustrate the flexibility of our local knockoff filter method.
Recall the high-level ideas of the procedure described in Section 2.5, also visualized schematically in Figure A2. For each pair $(j,\ell)$, we extract the scores $(T_{j,\ell},\tilde{T}_{j,\ell})$ from a local predictive model $\hat{\psi}^{(j,\ell)}$ trained using information from the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ and the local dataset $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$, which includes only observations $(X_{j},\tilde{X}_{j},Y,Z)$ for individuals $i\in[n]$ with $\nu_{j}(Z^{i})=\ell$. We now outline this implementation in more detail.
Figure A2: Schematic of the Adaptive Local Knockoff Filter. The relevant variables and associated statistics for different local environments are shown in different colors. The original variables are shown in darker shades and the knockoffs in lighter shades.
First, a sparse generalized linear model (e.g., the lasso) $\hat{\psi}^{(0)}$ is trained to predict $\mathbf{Y}$ given the data in $\tilde{\mathcal{D}}(\mathbf{V})$, tuning the regularization via cross-validation. The absolute values of the (scaled) regression coefficients from $\hat{\psi}^{(0)}$ will serve as “prior” importance measures, $\smash{T_{j}^{\text{prior}}}$ and $\smash{\tilde{T}_{j}^{\text{prior}}}$, for all variables and knockoffs indexed by $j\in[p]$. These are combined pairwise into a weight $\pi_{j}$ for each $j$; for example, as $\smash{\pi_{j}=\zeta(T^{\text{prior}}_{j}+\tilde{T}^{\text{prior}}_{j})}$, where $\zeta:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ is a positive and decreasing function such as $\zeta(t)=1/(0.05+t)$. Larger values of $\pi_{j}$ suggest the $j$-th variable is more likely to have a significant effect among some individuals. Similar weights can also be defined for the covariates.
Then, separately for each pair $(j,\ell)$, a local model $\hat{\psi}^{(j,\ell)}$ is trained with $\mathbf{Y}^{(j,\ell)}$ as the outcome and $[\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V}_{-j})},\mathbf{Z}]^{(j,\ell)}$ as the predictor matrix, where $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V}_{-j})}$ is defined as a narrower version of the cloaked data set $\tilde{\mathcal{D}}(\mathbf{V})$, in which the $j$-th and $(p+j)$-th columns are omitted. Omitting the $j$-th and $(p+j)$-th columns of $[\mathbf{X},\tilde{\mathbf{X}}]$ is intuitive since the true variable $\mathbf{X}_{j}$ and its knockoff $\tilde{\mathbf{X}}_{j}$ are already included in this local model.
To leverage any additional (non-local) information that may be contained in the cloaked data set $\tilde{\mathcal{D}}(\mathbf{V})$, each sparse local model $\hat{\psi}^{(j,\ell)}$ is trained with feature-specific regularization that depends on the prior weights $\pi$ extracted from $\hat{\psi}^{(0)}$. Specifically, for each $l\in[p]$, the penalty for the $l$-th and $(l+p)$-th variables in $\hat{\psi}^{(j,\ell)}$ is $\smash{\lambda^{(j,\ell)}_{l}=\lambda^{(j,\ell)}(1-\xi^{(j,\ell)})+\xi^{(j,\ell)}\pi_{l}}$, where $\lambda^{(j,\ell)}>0$ and $\xi^{(j,\ell)}\in[0,1]$ are two hyper-parameters to be tuned by cross-validation.
If $\smash{\xi^{(j,\ell)}=0}$, this reduces to a local model that only depends on the data in subgroup $(j,\ell)$. But, in general, larger values of $\xi^{(j,\ell)}$ can make our approach more powerful, as we gather strength from the data in all subgroups. In fact, null variables will tend to receive smaller values of $\pi$ and will thus be less likely to be incorrectly selected by the final model, thereby reducing the noise in the test statistics. Of course, the weights $\pi$ may not always be informative, hence why $\xi^{(j,\ell)}$ is tuned by cross-validation. For that purpose, an expensive two-dimensional grid search can be avoided by tuning first $\lambda^{(j,\ell)}>0$ and then $\xi^{(j,\ell)}$.
Finally, the importance scores for the $j$-th variable and knockoff in subgroup $\ell$ are respectively defined as the absolute values of the (scaled) regression coefficients for $X_{j}$ and $\tilde{X}_{j}$ in the local model $\hat{\psi}^{(j,\ell)}$. This procedure is summarized in Algorithm 3. The following result states that this leads to importance scores satisfying the equivariance property in (9).
Input: Data $\mathcal{D}=([\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z})$; Cloaked data set $\tilde{\mathcal{D}}(\mathbf{V})=([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z})$, Partition function $\nu=(\nu_{1},\ldots,\nu_{p})$;
Positive and decreasing function $\zeta:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$;
Finite lists of possible hyper-parameter values $\Lambda\subseteq\mathbb{R}_{+}$ and $\Xi\subseteq\mathbb{R}_{+}$;
1
2Fit a sparse generalized linear model $\hat{\psi}^{(0)}$ to predict $Y$, using the data in $\tilde{\mathcal{D}}(\mathbf{V})$;
3 for $j\in[p]$ do
4       Define $\smash{T_{j}^{\text{prior}}}$ and $\smash{\tilde{T}_{j}^{\text{prior}}}$ as the absolute values of the (scaled) regression coefficients in $\hat{\psi}^{(0)}$ for the $j$-th and $(j+p)$-th variables in $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}$, respectively;
5       Compute $\pi_{j}=\zeta(T^{\text{prior}}_{j}+\tilde{T}^{\text{prior}}_{j})$.
6Compute the partition width $L=\sum_{j=1}^{p}L_{j}$, where $L_{j}=\max_{z\in\mathcal{Z}}\nu_{j}(z)$;
7 for $j\in[p]$ do
8       for $\ell\in[L_{j}]$ do
9             Define the local data set $\tilde{\mathcal{D}}^{(j,\ell)}(\mathbf{V})=([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V}_{-j})},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$;
10             for $(\lambda,\xi)\in\Lambda\times\Xi$  do
11                   For each $l\in[p]$, define the weight $\smash{\lambda_{l}=\lambda(1-\xi)+\xi\pi_{l}}$;
12                   Fit a sparse generalized linear model $\hat{\psi}^{(j,\ell)}(\lambda,\xi)$ to predict $Y$ using the data in a cross-validated subset of $\tilde{\mathcal{D}}^{(j,\ell)}(\mathbf{V})$, with weights $(\lambda_{1},\ldots,\lambda_{p},\lambda_{1},\ldots,\lambda_{p},0,\ldots,0)$ for $([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V}_{-j})},\mathbf{Z})$;
13
14      Select the hyper-parameters $(\hat{\lambda}^{(j,\ell)},\hat{\xi}^{(j,\ell)})$ that minimize the cross-validated test prediction error for $Y$ based on the data in $\tilde{\mathcal{D}}^{(j,\ell)}(\mathbf{V})$;
15       For each $l\in[p]$, define the weight $\smash{\hat{\lambda}^{(j,\ell)}_{l}=\hat{\lambda}^{(j,\ell)}(1-\hat{\xi}^{(j,\ell)})+\hat{\xi}^{(j,\ell)}\pi_{l}}$;
16       Fit a sparse generalized linear model $\hat{\psi}^{(j,\ell)}$ for $Y$ using all data in $\tilde{\mathcal{D}}^{(j,\ell)}(\mathbf{V})$, with weights $(\hat{\lambda}^{(j,\ell)}_{1},\ldots,\hat{\lambda}^{(j,\ell)}_{p},\hat{\lambda}^{(j,\ell)}_{1},\ldots,\hat{\lambda}^{(j,\ell)}_{p},0,\ldots,0)$ for $([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V}_{-j})},\mathbf{Z})$;
17       Define $\smash{T_{j,\ell}}$ and $\smash{\tilde{T}_{j,\ell}}$ as the absolute values of the (scaled) regression coefficients in $\hat{\psi}^{(j,\ell)}$ for $X_{j}$ and $\tilde{X}_{j}$, respectively;
18
Output: List of importance scores $(T_{j,\ell},\smash{\tilde{T}_{j,\ell}})$ for all $j\in[p]$ and $\ell\in[L_{j}]$.
Algorithm 3 Computing powerful lasso-based local importance scores
Proposition A2.
The local importance scores $[\mathbf{T},\tilde{\mathbf{T}}]$ output by Algorithm 3 satisfy the equivariance property defined in (9).
Proof.
This is an immediate corollary of Proposition A3, stated in Appendix A3. ∎
A natural extension of this implementation would involve estimating different prior weights $\pi$ in different local environments, as sketched by Figure A3. This more sophisticated implementation still satisfies (9) as long as the models utilized to compute the priors only look at the cloaked data in $\tilde{\mathcal{D}}(\mathbf{V})$.
Figure A3: Schematic of Adaptive Local Knockoff Filter, focusing on the computation of the statistics for the first (blue) and fifth (green) local environments. The test statistics are based on a simple empirical prior for the two variables that is learnt from the cloaked data set with randomly swapped knockoffs. Other details are as in Figure A2.
A2.2 Learning Local Environments Using Interaction Models
This section describes a practical approach for learning knockoff-invariant partition functions that allow testing informative local conditional hypotheses with our method. This approach is based on parametric regression models with interaction terms.
Suppose that we suspect some of the variables denoted as $X$ may interact with certain covariates $Z$, akin to the toy example from Section 2.3. For simplicity, we focus on binary covariates here, although one could also similarly handle more general categorical covariates, or even continuous-valued covariates, through standard discretization techniques.
In this scenario, the analyst first generates a matrix $\mathbf{V}$ of i.i.d. Bernoulli random variables, independent of the data, and assembles $\tilde{\mathcal{D}}(\mathbf{V})$. Then, the analyst proceeds by fitting a sparse generalized linear model (i.e., the lasso) to predict $\mathbf{Y}$ given $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$ and $\smash{\mathbf{Z}}$, including all possible pairwise interactions between $\mathbf{Z}$ and $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$. Several algorithms exist for fitting such models (Bien et al., 2013; Lim and Hastie, 2015; Tibshirani and Friedman, 2020), making this approach practical and effective.
After tuning the lasso regularization via cross-validation, let $\hat{\beta}_{j}$ and $\hat{\beta}_{j+p}$ indicate the estimated main effects for the (scaled) variables in the $j$-th and $(j+p)$-th columns of $[\mathbf{X},\widetilde{\mathbf{X}}]_{\text{swap}(\mathbf{V})}$, respectively. Since $X_{j}$ and $\widetilde{X}_{j}$ are cloaked, in this model there is no meaningful distinction between the original variables and the knockoffs. Let also $\hat{\gamma}_{l,j}$ and $\hat{\gamma}_{l,j+p}$ indicate the corresponding interaction coefficients involving $Z_{l}$ and the variables in the $j$-th and $(j+p)$-th columns of $[\mathbf{X},\widetilde{\mathbf{X}}]_{\text{swap}(\mathbf{V})}$, respectively, for all $l\in[m]$ and $j\in[p]$.
Given an upper bound $G_{\max}$ on the number of interactions per variable (e.g., $G_{\max}=2$), for all $j\in[p]$ let ^$\smash{\hat{\mathcal{I}}_{j}\subset[m]}$ indicate the subset of $G_{\max}$ covariates with the strongest (nonzero) interactions involving the randomly swapped versions of $X_{j}$ or $\tilde{X}_{j}$. Here, we understand that any ties are broken at random and |^$\smash{|\hat{\mathcal{I}}_{j}|<G_{\max}}$ if the number of covariates with nonzero interaction coefficients is too small. In other words, we compute
^$\hat{\mathcal{I}}_{j}:=\big{\{}l\in[m]:|\hat{\gamma}_{l,j}|+|\hat{\gamma}_{l,j+p}|>0,\sum_{l^{\prime}\neq l}\mathbbm{1}\left[|\hat{\gamma}_{l,j}|+|\hat{\gamma}_{l,j+p}|>|\hat{\gamma}_{l^{\prime},j}|+|\hat{\gamma}_{l^{\prime},j+p}|\right]>m-G_{\max}\big{\}}.$
This model thus links each pair $(X_{j},\tilde{X}_{j})$ to at most $G_{\max}$ covariates. The corresponding functions $\smash{\hat{\nu}_{j}}$ take values in $\smash{\{1,\ldots,2^{|\hat{\mathcal{I}}_{j}|}\}}$, which indexes all possible configurations of the covariates in ^$\hat{\mathcal{I}}_{j}$. Therefore, we may write $|\hat{L}_{j}|=2^{|\hat{\mathcal{I}}_{j}|}$ for each $j\in[p]$.
For example, in the special case where ^$\hat{\mathcal{I}}_{j}=\emptyset$, meaning that no interaction effects are selected for the pair $(X_{j},\tilde{X}_{j})$, this procedure yields $\nu_{j}(z)=1$ for all possible $z\in\mathbb{R}^{m}$. In other words, in this case, the partition function $\hat{\nu}_{j}$ defines a single local environment for the $j$-th variable, comprising the full sample. By contrast, if |^$|\hat{\mathcal{I}}_{j}|=\{l\}$ for some $l\in[m]$, the partition function $\hat{\nu}_{j}$ will define two local environments for the $j$-th variable: one corresponding to $Z_{l}=0$ and the other to $Z_{l}=1$.
Table A1 helps visualize this approach by highlighting a thought experiment involving 3 explanatory variables and 3 binary covariates. In this toy example, we imagine that the lasso selects $Z_{1},Z_{2}$ as potentially interacting with $X_{1}$, and $Z_{3}$ for $X_{2}$. No interactions with $X_{3}$ are detected. Thus, the population is partitioned into 4 subgroups for $X_{1}$, 2 subgroups for $X_{2}$, and 1 trivial subgroup for $X_{3}$, yielding a total of 7 data-driven hypotheses.
Table A1: Subgroups selected by fitting a lasso model with interactions, and corresponding interpretations of the possible discoveries obtained with our method, in a toy example with 3 variables and 3 binary covariates. The term “influences” is utilized loosely here to indicate a (possibly non-causal) significant conditional association.
Variable Covariates Partition Interpretation of the findings (upon rejection)
Label Definition
$X_{1}$$Z_{1},Z_{2}$} 1 $Z_{1}=0,Z_{2}=0$ $X_{1}$ influences $Y$ if $Z_{1}=0$ and $Z_{2}=0$
2 $Z_{1}=0,Z_{2}=1$ $X_{1}$ influences $Y$ if $Z_{1}=0$ and $Z_{2}=1$
3 $Z_{1}=1,Z_{2}=0$ $X_{1}$ influences $Y$ if $Z_{1}=1$ and $Z_{2}=0$
4 $Z_{1}=1,Z_{2}=1$ $X_{1}$ influences $Y$ if $Z_{1}=1$ and $Z_{2}=1$
$X_{2}$$Z_{3}$} 1 $Z_{3}=0$ $X_{2}$ influences $Y$ if $Z_{3}=0$
2 $Z_{3}=1$ $X_{2}$ influences $Y$ if $Z_{3}=1$
$X_{3}$emptyset 1 All individuals $X_{3}$ influences $Y$
Note that the parameter $G_{\max}$ generally controls an important trade-off between the power of our method and the interpretability of the findings. A larger value of $G_{\max}$ tends to lead to more specific hypotheses corresponding to smaller subgroups, but it also makes it more difficult to reject those hypotheses by reducing the effective sample size available to compute the local importance scores by a factor of up to $2^{G_{\max}}$, in the case of binary covariates. Concretely, all demonstrations presented in this paper utilize $G_{\max}\in\{1,2\}$, but the optimal value of $G_{\max}$ may generally depend on the data.
A2.3 Implementing aLKF for GWAS Data
One of the motivations behind the methods developed in this paper is to study how genetic variants influencing a phenotype may vary across environments, often referred to as “gene-environment interactions”. Consequently, it is important to describe a version of the aLKF that can be applied effectively and without excessive computational costs to the analysis of genetic data. Key characteristics to consider are that the number of potential explanatory variables, $p$, typically ranges from hundreds of thousands to millions, while the number of measured covariates, $m$, is much smaller, often on the order of tens.
These characteristics make it impractical to implement our method as described in the previous sections, creating a need for a specialized implementation. This implementation departs from the approach described so far in three main aspects.
  • Pre-screening variables. A lasso model fitted to the cloaked data identifies a subset of promising variables ($p_{\text{screen}}\ll p$) by selecting variables with non-zero lasso coefficients for either version of the cloaked original and knockoff SNPs.
  • Learning local environments. A lasso model with interaction terms is applied to the pre-screened data to define data-driven partition functions. To prioritize global associations and avoid unnecessary sample splits, the penalty for interaction terms is scaled by a hyperparameter $c\in(0,1]$ (e.g., $c=0.25$).
  • Batch computation of importance scores. Variables with identical partition functions are grouped into batches. A single predictive model is fitted per batch, reducing the number of models needed, especially when $m\ll p_{\text{screen}}$.
Algorithms 4 and 5 summarize this specialized implementation of our method, with additional details provided in Appendices A2.3.1A2.3.3. Furthermore, Appendix A2.3.4 describes how our method can be applied after partitioning the genome into blocks of SNPs. This accounts for strong linkage disequilibrium (LD) among nearby variants on the same chromosome, enabling a multi-resolution analysis (Sesia et al., 2020, 2021).
Input: Pre-screened variables $\mathbf{X}$ and knockoffs $\tilde{\mathbf{X}}$, outcomes $\mathbf{Y}$, covariates $\mathbf{Z}$;
Random noise matrix $\mathbf{V}$; Partition function $\nu=(\nu_{1},\ldots,\nu_{p_{\text{screen}}})$;
Nominal FDR level $\alpha\in(0,1)$;
1
2Compute the partition width $L=\sum_{j=1}^{p_{\text{screen}}}L_{j}$, where $L_{j}=\max_{z\in\mathcal{Z}}\nu_{j}(z)$;
3 Define local hypotheses $\mathcal{H}_{j,\ell}(\nu)$ as in (5), for $j\in[p_{\text{screen}}]$ and $\ell\in[L_{j}]$;
4 Partition all $j\in[p_{\text{screen}}]$ into batches $\mathcal{Q}$ such that $\hat{\nu}_{j}(\mathbf{Z})=\hat{\nu}_{k}(\mathbf{Z})$ $\forall$ $j,k\in Q$;
5 Assemble the cloaked data matrix $\tilde{\mathcal{D}}(\mathbf{V})=([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z})$;
6 for $Q\in\mathcal{Q}$ do
7       for $\ell\in[L_{Q}]$ do
8             Fit a lasso model of $\mathbf{Y}^{(Q,\ell)}$ on $[[\mathbf{X}_{j},\tilde{\mathbf{X}}_{j}]_{j\in Q},[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V},-Q)},\mathbf{Z}]^{(Q,\ell)}$;
9             for $j\in Q$ do
10                   Define $T_{j,\ell}=|\hat{\beta}_{j,\ell}|$ and $\tilde{T}_{j,\ell}=|\hat{\beta}_{j+|Q|,\ell}|$, where $\hat{\beta}$ are the lasso coefficients;
11                   Compute the test statistic $W_{j,\ell}=T_{j,\ell}-\tilde{T}_{j,\ell}$ for all $j\in\mathcal{Q}$;
12
13Vectorize the test statistics $W_{j,\ell}$ for all $j\in[p_{\text{screen}}],\ell\in[L_{j}]$ into $\mathbf{W}$;
14 Apply Selective SeqStep+ (Barber and Candès, 2015) at level $\alpha$ to the vector $\mathbf{W}$;
15
Output: List of rejected hypotheses ^$\hat{\mathcal{R}}\subseteq\{\mathcal{H}_{j,\ell}(\nu)\}_{j\in[p_{\text{screen}}],\ell\in[L_{j}]}$.
Algorithm 4 Local Knockoff Filter for GWAS data, with fixed hypotheses
Input: Knockoff-augmented data set $\mathcal{D}=([\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z})$; FDR level $\alpha\in(0,1)$;
1 Randomly generate $\mathbf{V}\in\{0,1\}^{n\times p}$, a matrix of i.i.d. Bernoulli noise;
2 Assemble the cloaked data matrix $\tilde{\mathcal{D}}(\mathbf{V})=([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z})$;
3 Obtain pre-screened variants $[\mathbf{X},\mathbf{\widetilde{X}}]_{\text{(swap({V}), screen)}}$ by fitting a lasso model for the outcome $\mathbf{Y}$ given $([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Z})$; see Appendix A2.3.1;
4 Learn a partition function $\hat{\nu}$, with width $\hat{L}=\sum_{j\in[p_{\text{screen}}]}\hat{L}_{j}$, by fitting a lasso model of $\mathbf{Y}$ on $[[\mathbf{X},\mathbf{\widetilde{X}}]_{\text{(swap({V}), screen)}},\mathbf{Z}]$, including interactions; see Appendix A2.3.2;
5 Apply Algorithm 4 using the partition function $\hat{\nu}$;
Output: List of rejected hypotheses ^$\hat{\mathcal{R}}\subseteq\{\mathcal{H}_{j,\ell}\}_{j\in[p_{\text{screen}}],\ell\in[\hat{L}_{j}]}$.
Algorithm 5 Adaptive Local Knockoff Filter for GWAS data
A2.3.1 Pre-Screening of Promising Variables Using Cloaked Data
To reduce the number of local hypotheses that need to be considered, while maintaining control over the FDR for the selected hypotheses, we begin by identifying a smaller set of promising explanatory variables, $p_{\text{screen}}<p$ (e.g., with $p_{\text{screen}}$ on the order of tens of thousands), through a pre-screening step based on the cloaked data in $\tilde{\mathcal{D}}(\mathbf{V})$. The goal of this step is to reduce the dimensionality of the problem while retaining variables that are likely to be relevant.
To pre-screen relevant variables, we fit a lasso model of $\mathbf{Y}$ on $[[\mathbf{X},\widetilde{\mathbf{X}}]_{\text{swap}(\mathbf{V})},\mathbf{Z}]$, tuning the regularization parameter via cross validation. Note that computationally efficient software packages (Privé et al., 2018, 2019) make this task feasible even for large-scale GWAS data (Sesia et al., 2020, 2021), as this model only needs to be fitted once and does not include interaction terms.
Let $\hat{\beta}_{j}$ and $\hat{\beta}_{j+p}$ indicate the estimated lasso coefficients for the (scaled) variables in the $j$-th and $(j+p)$-th columns of $[\mathbf{X},\widetilde{\mathbf{X}}]_{\text{swap}(\mathbf{V})}$, respectively. Since $X_{j}$ and $\widetilde{X}_{j}$ are cloaked, in this model there is no meaningful distinction between the original variables and the knockoffs. Therefore, the pre-screening selects both columns containing the randomly swapped values of $X_{j}$ and $\widetilde{X}_{j+p}$, for all $j\in[p]$ such that $|\hat{\beta}_{j}|+|\hat{\beta}_{j+p}|>0$.
To simplify the notation in the following, we will assume, by convention and without loss of generality, that the indices $j\in[p]$ corresponding to the $p_{\text{screen}}$ original variables that passed the screening step are given by $[p_{\text{screen}}]=\{1,\ldots,p_{\text{screen}}\}$. Furthermore, we will denote by $[\mathbf{X},\widetilde{\mathbf{X}}]_{\text{swap}(\mathbf{V}),\text{screen}}$ the data matrix containing the $p_{\text{screen}}$ (cloaked) pairs of variables and knockoffs that passed the screening step.
A2.3.2 Learning Local Environments Using Pre-Screened Data
Fitting a lasso model with interaction terms.
The second step of our method focuses on learning a potentially interesting partition function $\hat{\nu}$ by fitting a lasso model with interaction terms, similar to the approach described in Appendix A2.2. Again, for simplicity we are assuming here that the covariates are binary, as in Appendix A2.2. The key difference is that, instead of using the full cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})=(\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z}})$, now we use its smaller pre-screened version, $\tilde{\mathcal{D}}_{\text{screen}}(\mathbf{V})=(\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V}),\text{screen}},\mathbf{Y},\mathbf{Z}})$.
In particular, we fit a lasso model to predict $\mathbf{Y}$ given $[[\mathbf{X},\mathbf{\widetilde{X}}]_{\text{(swap({V}), screen)}},\mathbf{Z}]$, after augmenting the data matrix with all possible pairwise interactions between $\mathbf{Z}$ and $[\mathbf{X},\mathbf{\widetilde{X}}]_{\text{(swap({V}), screen)}}$, and tuning the regularization parameter via cross validation. Again, let $\hat{\beta}_{j}$ and $\hat{\beta}_{j+p}$ indicate the estimated main effects for the $j$-th and $(p+j)$-th variables in $[\mathbf{X},\mathbf{\widetilde{X}}]_{\text{(swap({V}), screen)}}$, for any $j\in[p_{\text{screen}}]$. Further, let $\hat{\gamma}_{l,j}$ and $\hat{\gamma}_{l,j+p}$ indicate the corresponding interaction coefficients involving $Z_{l}$, for all $l\in[m]$ and $j\in[p_{\text{screen}}]$.
Finally, we define the data-driven partition function $\hat{\nu}$ with the same approach described in Appendix A2.2. In particular, the interaction effect for $Z_{l}$ with $l\in[m]$ is selected whenever $|\hat{\gamma}_{l,j}|+|\hat{\gamma}_{l,j+p}|>0$ and the main effect if $|\hat{\gamma}_{l,j}|+|\hat{\gamma}_{l,j+p}|=0$ but $|\hat{\beta}_{j}|+|\hat{\beta}_{j+p}|>0$. As in Appendix A2.2, we consider only the $G_{\max}$ covariates with the strongest (nonzero) interactions involving $X_{j}$ or $\tilde{X}_{j}$, where $G_{\max}$ is a fixed hyper-parameter.
Encouraging the selection of coarser partitions.
Large-scale GWAS data often exhibit two additional characteristics that must be considered when applying the aLKF method in order to maximize power. First, in the analysis of polygenic traits, there are typically many discoveries to be made. The global knockoff filter of Candès et al. (2018) already demonstrates high power, often identifying hundreds or thousands of significant associations between genetic variants and the phenotype (Sesia et al., 2020, 2021). Second, not all genetic variants are likely to have subgroup-specific effects. For many significant SNPs, it may be sufficient to detect associations at the population level. Only a relatively small fraction of global associations are expected to correspond to effects specific to particular subgroups. In other words, not all significant genetic effects are expected to involve interactions with covariates.
These considerations highlight the need to ensure that the aLKF method does not unnecessarily reduce power to detect global associations relative to existing methods. A key concern is that if the lasso model with interactions described above may select too many spurious interaction terms, the data-driven partition function $\hat{\nu}$ may unnecessarily split the sample into overly small subgroups. This would reduce the sample size available for computing local importance scores and potentially diminish power.
To address this issue, the lasso model with interactions must be calibrated more conservatively by increasing the practical threshold for selecting interaction terms, thus creating an incentive to select coarser sample partitions. In practice, this is implemented by adjusting the lasso penalty applied to main effects. Specifically, the penalty for main effects in the interaction model described above is multiplied by a factor $c\in(0,1]$, with smaller values of $c$ effectively increasing the cost of selecting interaction terms. While the choice of $c$ involves some arbitrariness, our numerical experiments have shown that $c=0.25$—which emphasizes global effects—works well in practice across a variety of scenarios, involving synthetic data with different proportions of true local vs. global effects.
A2.3.3 Batch Computation of Local Importance Scores
Recall that typical GWAS datasets are characterized by a very large number, $p$, of explanatory variables and a much smaller number, $m$, of relevant covariates. Even after pre-screening, we generally expect $p_{\text{screen}}$ to remain much larger than $m$. This imbalance creates a natural opportunity to streamline the computation of the importance scores.
Rather than re-fitting a local predictive model for each variable $j\in[p_{\text{screen}}]$ and subgroup $\ell\in\hat{L}_{j}$, we can perform “batch computation”: the key intuition is that it is possible to utilize the same local predictive models for different variables assigned to identical partition functions. When $m$ is much smaller than $p_{\text{screen}}$, this shortcut significantly reduces the number of local models that need to be fitted compared to Algorithm 3.
To illustrate this process, consider a simplified setting with $p_{\text{screen}}=4$ explanatory variables, post-screening, and a single binary covariate, $Z_{1}\in\{0,1\}$. Suppose the detected local environments are as described in Table A2 below. Based on the subgroups identified in Table A2, Algorithm 1 would require fitting six distinct predictive models. In contrast, the batch processing shortcut described below reduces this to just three local models: one for the entire sample, one for the subgroup defined by $Z_{1}=0$ and one for the subgroup defined by $Z_{1}=1$.
Table A2: Partition functions learnt by fitting a lasso model with interactions in an imaginary toy example with 4 variables and a single binary covariate.
Variable Covariates Partition
Label Definition
$X_{1}$emptyset 1 All individuals
$X_{2}$$Z_{1}$} 1 $Z_{1}=0$
2 $Z_{1}=1$
$X_{3}$$Z_{1}$} 1 $Z_{1}=0$
2 $Z_{1}=1$
$X_{4}$emptyset 1 All individuals
Our “batch computation” shortcut can be generally described as follows.
Defining the batches.
First, we partition $[p_{\text{screen}}]$ into disjoint “batches” $\mathcal{Q}$, such that $[p_{\text{screen}}]=\cup_{Q\in\mathcal{Q}}Q$ and all variables within each batch have the same partition function: $\hat{\nu}_{j}(\mathbf{Z})=\hat{\nu}_{k}(\mathbf{Z})$ for all $j,k\in Q$, for all $Q\in\mathcal{Q}$. Let this joint partition be denoted by $\hat{\nu}_{Q}$, with $\hat{L}_{Q}(\nu)$ representing the number of disjoint batchs induced by $\nu_{Q}$, and let the associated covariates be denoted by $\mathbf{Z}_{Q}\subseteq\mathbf{Z}$.
In the toy example of Table A2, the partition function shared by $X_{2}$ and $X_{3}$ divides the sample into two subgroups corresponding to $Z_{1}=0$ and $Z_{1}=1$, while the partition function for $X_{1}$ and $X_{4}$ does not partition the sample. Based on this, we would assign $X_{1}$ and $X_{4}$ to the first batch and $X_{2}$ and $X_{3}$ to the second batch.
Computing importance scores via batch-local predictive models.
Separately for each batch $Q\in\mathcal{Q}$ and subgroup $\ell\in\hat{L}_{Q}$, we fit a “batch-local” predictive model, denoted as $\hat{\psi}^{(Q,\ell)}$, for $Y$ given $(X,\tilde{X},Z)$ as follows. This model can be trained using information from the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ and a smaller batch-local dataset $\mathcal{D}^{(Q,\ell)}=([\mathbf{X}_{j},\tilde{\mathbf{X}}_{j}]_{j\in Q},\mathbf{Y},\mathbf{Z})^{(Q,\ell)}$, which includes only observations $(X_{j},\tilde{X}_{j},Y,Z)$ for all $j\in Q$ and all individuals $i\in[n]$ with $\nu_{j}(Z^{i})=\ell$. Concretely, we implement $\hat{\psi}^{(Q,\ell)}$ using the lasso, trained with $\mathbf{Y}^{(Q,\ell)}$ as the outcome and $[[\mathbf{X}_{j},\tilde{\mathbf{X}}_{j}]_{j\in Q},[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V},-Q)},\mathbf{Z}]^{(Q,\ell)}$ as the predictor matrix, where $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V},-Q)}$ denotes the matrix $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}$ after removing the columns indexed by $j$ and $p+j$ for all $j\in Q$. The sparsity parameter of this lasso model is tuned as usual, via cross-validation.
Then, to obtain local importance scores satisfying (9), it is sufficient to compute $(T_{j,\ell},\tilde{T}_{j,\ell})$, for each $j\in Q$, in such a way that swapping $X_{j}$ and $\tilde{X}_{j}$ in $\mathcal{D}^{(Q,\ell)}$ would result in $T_{j,\ell}$ being swapped with $\tilde{T}_{j,\ell}$; this is analogous to Proposition A3 in Appendix A3. Such equivariance is easily achieved by setting $T_{j,\ell}=|\hat{\beta}_{j,\ell}|$ and $\tilde{T}_{j,\ell}=|\hat{\beta}_{j+|Q|,\ell}|$, for each $j\in Q$, where $\hat{\beta}_{j,\ell}$ and $\hat{\beta}_{j+|Q|,\ell}$ respectively denote the (scaled) estimated regression coefficients of variable $j$ and its knockoff within the batch-local lasso model $\hat{\psi}^{(Q,\ell)}$.
Note that, unlike the procedure in Appendix A2.1, when computing local scores within a batch $(Q,\ell)$, this approach does not utilize any additional (non-local) information from the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ beyond what is captured by $[[\mathbf{X}_{j},\tilde{\mathbf{X}}_{j}]_{j\in Q},[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V},-Q)},\mathbf{Z}]^{(Q,\ell)}$. Specifically, the “prior” regularization weights described in Appendix A2.1 are not used, effectively setting the hyper-parameter $\smash{\xi^{(Q,\ell)}=0}$. The motivation for this choice is two-fold. First, since the number of covariates $m$ is expected to be very small compared to both the number $p$ of variables and the sample size $n$, the effective sample size of $[[\mathbf{X}_{j},\tilde{\mathbf{X}}_{j}]_{j\in Q},[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V},-Q)},\mathbf{Z}]^{(Q,\ell)}$ is already relatively large. Second, having fewer hyper-parameters to tune via cross-validation is helpful to further reduce computational costs.
Finally, after evaluating all local importance scores, we proceed as described in Section 2.5, with test statistics $W_{j,\ell}=T_{j,\ell}-\tilde{T}_{j,\ell}$ for each $Q\in\mathcal{Q}$, $j\in Q$ and $\ell\in\hat{L}_{Q}$.
A2.3.4 Accounting for Strong Linkage Disequilibrium
Due to the presence of strong local dependencies among genetic variants on the same chromosome, a phenomenon known as linkage disequilibrium, it is often intrinsically very challenging to precisely identify individual variants that influence a complex trait. Instead, significant genetic associations are typically detected at the level of broader genetic loci, which include several SNPs in strong linkage disequilibrium.
To address this complication in GWAS data, Sesia et al. (2020) and Sesia et al. (2021) applied the global knockoff filter of Candès et al. (2018) at different levels of resolution. Specifically, they partitioned the genome into contiguous blocks of variants in strong linkage disequilibrium and tested for conditional independence with the phenotype at the block level, using these blocks as the units of inference rather than individual SNPs.
The same strategy can be seamlessly integrated with our aLKF method.
Starting with a pre-defined partition $\mathcal{B}$ of the genome into contiguous blocks of SNPs $b\in\mathcal{B}$ at the desired resolution, along with corresponding knockoff genotypes, as described in Sesia et al. (2020) and Sesia et al. (2021), we can apply our method as outlined so far, with four key modifications:
  1. 1.
    The data cloaking step using the random matrix $\mathbf{V}$ is performed at the block level, assigning the same values of $V_{ij}$ to all SNPs $j$ within a block $b\in\mathcal{B}$. This reflects the idea that the units of inference are now blocks of SNPs and not single SNPs.
  2. 2.
    Each block of SNPs $b$ is assigned a single partition function $\hat{\nu}_{b}$, rather than having distinct partition functions for individual SNPs.
  3. 3.
    While pre-screening nominally selects a subset of individual variants, since inference is then performed at the block level, the batch-local predictive models are fitted using not only the pre-screened variants but also all other variants in the same block.
  4. 4.
    For each combination of $b$ and $\ell\in\hat{L}_{b}$, a single pair of local scores $(T_{b,\ell},\tilde{T}_{b,\ell})$ is computed by summing the single-SNP scores corresponding to the individual SNPs within that block. Then, the corresponding test statistic is $W_{b,\ell}=T_{b,\ell}-\tilde{T}_{b,\ell}$.
Finally, the discoveries obtained with our method will be interpreted as highlighting blocks of SNPs displaying significant associations with the outcome within certain subgroups determined by specific covariate values.
Appendix A3 Mathematical Proofs
Proposition A3.
Consider a vector of local importance scores ${[\mathbf{T},\tilde{\mathbf{T}}]}$ which can be written as the output of a suitable (randomized) function $\bm{\tau}$ applied to $\mathcal{D}$ and $\tilde{\mathcal{D}}(\mathbf{V})$; that is, $[\mathbf{T},\tilde{\mathbf{T}}]=\bm{\tau}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)}=\big{[}\bm{t}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)},\tilde{\bm{t}}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)}\big{]}$, using the same notation as in (8). Suppose that, for any pair $(j,\ell)$, the scores $(T_{j,\ell},\smash{\tilde{T}_{j,\ell}})$ are a function of a local model $\hat{\psi}^{(j,\ell)}$, trained looking only at $\tilde{\mathcal{D}}(\mathbf{V})$ and $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$, and satisfy the local symmetry property defined in (10). Then, $[\mathbf{T},\tilde{\mathbf{T}}]$ must satisfy the equivariance property defined in (9); that is, for any $\mathcal{S}\subseteq[\smash{L}]$, whose elements uniquely identify pairs $(j,\ell)$ for $j\in[p]$ and $\smash{\ell\in[{L}_{j}]}$,
$\bm{\tau}\big{(}\mathcal{D}_{\mathrm{swap}(\mathcal{S})},\tilde{\mathcal{D}}(\mathbf{V})\big{)}=\big{[}\bm{t}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)},\tilde{\bm{t}}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)}\big{]}_{\mathrm{swap}(\mathcal{S})}.$
Proof of Proposition A3.
This proof is quite straightforward. Consider how ${[\mathbf{T},\tilde{\mathbf{T}}]}$ would change if the input data set $\mathcal{D}$ were replaced by $\mathcal{D}_{\mathrm{swap}(\mathcal{S})}$, while keeping the cloaked data set $\tilde{\mathcal{D}}(\mathbf{V})$ unchanged.
  • For any $(j,\ell)\notin\mathcal{S}$, the local model $\hat{\psi}^{(j,\ell)}$ is invariant because both $\tilde{\mathcal{D}}(\mathbf{V})$ and $\mathcal{D}^{(j,\ell)}$ are invariant, as explained above. Therefore, in this case $\smash{T_{j,\ell}}$ and $\smash{\tilde{T}_{j,\ell}}$ are invariant.
  • For any $(j,\ell)\in\mathcal{S}$, the local model $\hat{\psi}^{(j,\ell)}$ remains invariant up to swapping the importance score of $X_{j}$ with that of $\tilde{X}_{j}$, because $\tilde{\mathcal{D}}(\mathbf{V})$ is invariant while $X_{j}$ and $\tilde{X}_{j}$ are swapped in $\mathcal{D}^{(j,\ell)}$, by (10). Therefore, in this case $\smash{T_{j,\ell}}$ and $\smash{\tilde{T}_{j,\ell}}$ are swapped.
This concludes the proof that
$\bm{\tau}\big{(}\mathcal{D}_{\mathrm{swap}(\mathcal{S})},\tilde{\mathcal{D}}(\mathbf{V})\big{)}=\big{[}\bm{t}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)},\tilde{\bm{t}}\big{(}\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V})\big{)}\big{]}_{\mathrm{swap}(\mathcal{S})}.$
Proof of Theorem 1.
Let $\smash{\bar{\mathbf{U}}\in\{\pm 1\}^{L}}$ be a random vector with independent entries such that: $\smash{\bar{U}_{j,\ell}=\pm 1}$ with probability $1/2$ if $\smash{\mathcal{H}_{j,\ell}(\nu)}$ in (5) is true and ${\bar{U}_{j,\ell}=+1}$ otherwise, for all $j\in[p]$ and ${\ell\in[L_{j}]}$. To prove Theorem 1, it suffices to prove the flip-sign property
[…]
since from there the result follows immediately as in Candès et al. (2018).
To prove (A4), we begin by recalling some useful notation. For each $j\in[p]$, ${L}_{j}(\nu)=\max_{z\in\mathcal{Z}}\nu_{j}(z)\in\mathbb{N}$ represents the number of disjoint subgroups induced by $\nu_{j}$. Define $\mathcal{S}(\bar{\mathbf{U}}):=\{(j,\ell):j\in[p],\ell\in[L_{j}],\bar{U}_{j,\ell}=-1\}\subseteq[L]$, where $L=\sum_{j=1}^{p}{L}_{j}$. Recall also that $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathcal{S}(\bar{\mathbf{U}}))}}$ is defined as the matrix obtained from $\smash{[\mathbf{X},\tilde{\mathbf{X}}]}$ after swapping the sub-column $\smash{\mathbf{X}^{(\ell)}_{j}}$, which contains all observations of $X_{j}$ in the subgroup $g_{j,\ell}$, with the corresponding knockoffs, for all $(j,\ell)\in\mathcal{S}(\bar{\mathbf{U}})$.
To emphasize that $\mathbf{W}$ is a function of $\mathcal{D}$ and $\tilde{\mathcal{D}}(\mathbf{V})$, we write $\mathbf{W}=\mathbf{W}(\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V}))$. Specifically, recall that $\mathbf{W}$ is obtained by applying an anti-symmetric combination function to $[\mathbf{T},\tilde{\mathbf{T}}]$ in (8). Further, Proposition A3 implies the importance scores ${[\mathbf{T},\tilde{\mathbf{T}}]}$ computed by Algorithm 1 satisfy (9). Therefore, it follows from (9) that
$\mathbf{W}(\mathcal{D}_{\mathrm{swap}(\mathcal{S}(\bar{\mathbf{U}}))},\tilde{\mathcal{D}}(\mathbf{V}))=\bar{\mathbf{U}}\odot\mathbf{W}(\mathcal{D},\tilde{\mathcal{D}}(\mathbf{V}))$
almost surely with respect to $\bar{\mathbf{U}},\mathbf{V}$, and $\mathcal{D}$, where $\mathbf{W}(\mathcal{D}_{\mathrm{swap}(\mathcal{S}(\bar{\mathbf{U}}))},\tilde{\mathcal{D}}(\mathbf{V}))$ represents the imaginary vector of test statistics which would be obtained by applying our method to the modified data set $\mathcal{D}_{\mathrm{swap}(\mathcal{S}(\bar{\mathbf{U}}))}=([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathcal{S}(\bar{\mathbf{U}}))},\mathbf{Y},\mathbf{Z})$ instead of $\mathcal{D}=(\smash{[\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z}})$.
Finally, the proof of (A4) is completed by showing that it follows from Lemma A1 that
[…]
for any subset $S\subseteq[L]$ of true null hypotheses $\mathcal{H}_{j,\ell}(\nu)$ in (5), which in turn implies
[…]
To see why Lemma A1 implies (A5), note that we can define a matrix $\mathbf{U}\in\{\pm 1\}^{n\times p}$ such that, for each $i\in[n]$ and $j\in[p]$, $U_{i,j}=\bar{U}_{j,l}$ where $\ell=\nu_{j}(Z^{i})$. In particular, this means that $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(S(\bar{\mathbf{U}}))}=[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{U})}$. Further, it follows from the definition of $\bar{\mathbf{U}}$ and $\mathcal{H}_{j,\ell}(\nu)$ in (5) that almost surely $U_{i,j}=1$ if $\mathcal{H}^{i}_{j}$ in (4) is not true. Therefore, we can directly apply Lemma A1 with this choice of $\mathbf{U}$ and obtain (A5).
Lemma A1.
Let $\mathbf{V}\in\{\pm 1\}^{n\times p}$ be i.i.d. Bernoulli random variables, independent of everything else. Under the model in Equation (3), consider any fixed matrix $\mathbf{U}\in\{\pm 1\}^{n\times p}$ such that, for each $i\in[n]$ and $j\in[p]$, $U_{i,j}=1$ if $\mathcal{H}^{i}_{j}$ in (4) is not true. Let $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{U})}\in\mathbb{R}^{n\times 2p}$ be the concatenation of $\mathbf{X}$ and $\tilde{\mathbf{X}}$, with the $i$-th observation of $X_{j}$ swapped with its knockoff if and only if $U_{ij}=-1$. Then,
[…]
which also implies
[…]
Proof of Lemma A1.
By independence, it suffices to prove this for a single row of the data matrices, which corresponds to an individual $i\in[n]$ with a true conditional distribution $p^{(i)}_{Y\mid X,Y}$. Keeping the individual’s index $i$ implicit to simplify the notation, we aim to prove
[…]
where $u\in\{-1,1\}^{p}$ satisfies $u_{j}=1$ for all variables $j\in[p]$ that are true nulls for the $i$-th individual, and $v$ denotes the $i$-th row of $\mathbf{V}$.
By construction of the knockoffs $\tilde{X}$, we know that
[…]
Combining (A6) and (A7) yields:
[…]
Now, recall that $Y\perp\!\!\!\perp\tilde{X}\mid X,Z$. Further, $Y\perp\!\!\!\perp X_{s}\mid X_{-s},Z$ for all $X_{s}$ swapped by $s$ because the latter only involves null variables. Therefore, by the same argument as in the proof of Lemma 3.2 in Candès et al. (2018),
[…]
Finally, combining (A8) with (A9) gives the desired result. ∎
Proof of Theorem 2.
[…]
Recall also that the test statistics $\mathbf{W}$ are determined by applying a (possibly randomized) function $\bm{\tau}$ to the knockoff-augmented data set $\mathcal{D}=([\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z})$ and $\tilde{\mathcal{D}}(\mathbf{V})$. Since the possible source of randomness in $\bm{\tau}$ is also independent of everything else, we may ignore it without loss of generality, to simplify the notation as much as possible.
Define $\mathcal{S}(\bar{\mathbf{U}}):=\{(j,\ell):j\in[p],\ell\in[L_{j}],\bar{U}_{j,\ell}=-1\}\subseteq[L]$, where $\hat{L}=\sum_{j=1}^{p}\hat{L}_{j}$. Recall also that $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathcal{S}(\bar{\mathbf{U}}))}}$ is defined as the matrix obtained from $\smash{[\mathbf{X},\tilde{\mathbf{X}}]}$ after swapping the sub-column $\smash{\mathbf{X}^{(\ell)}_{j}}$, which contains all observations of $X_{j}$ in the subgroup $g_{j,\ell}$, with the corresponding knockoffs, for all $(j,\ell)\in\mathcal{S}(\bar{\mathbf{U}})$. With this notation, it follows from (9) that
$\mathbf{W}([\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathcal{S}(\bar{\mathbf{U}}))},\mathbf{Y},\mathbf{Z},\tilde{\mathcal{D}}(\mathbf{V}))=\bar{\mathbf{U}}\odot\mathbf{W}([\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z},\tilde{\mathcal{D}}(\mathbf{V})).$ (A10)
Next, note that we can equivalently write (almost surely) $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\bar{\mathbf{U}})}=[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{U})}$, where $\mathbf{U}\in\{\pm 1\}^{n\times p}$ is a matrix defined such that, for each $i\in[n]$ and $j\in[p]$, $U_{i,j}=\bar{U}_{j,\ell}$ for $\ell=\hat{\nu}_{j}(Z_{i})$. In particular, this definition of $\mathbf{U}$ implies that: (1) $\mathbf{U}$ may be treated as fixed conditional on $\tilde{\mathcal{D}}(\mathbf{V})$, similar to $\bar{\mathbf{U}}$; and (2) $U_{i,j}=1$ whenever $\mathcal{H}^{i}_{j}$ in (4) is not true. Therefore, we can apply Lemma A1, obtaining that, conditional on $\tilde{\mathcal{D}}(\mathbf{V})$ and $\mathbf{U},\bar{\mathbf{U}}$,
[…]
and thus, conditional on $\tilde{\mathcal{D}}(\mathbf{V})$ and $\mathbf{U},\bar{\mathbf{U}}$,
[…]
Finally, combining Equations (A10) and (A11) leads to the desired result:
[…]
Appendix A4 Numerical Experiments
A4.1 Additional Details on the Experiments from Section 3.1
We provide here more details on the joint distribution of the variables utilized in the numerical Section 3.1, and on the corresponding construction of knockoffs.
The first 20 covariates are independently sampled from a Bernoulli(0.5) distribution:
[…]
The remaining 60 covariates are generated by applying the normal CDF transform to a Gaussian random vector with an order-one autoregressive covariance matrix and correlation parameter 0.5. That is,
$Z_{j}=\Phi(Z_{j}^{\text{latent}}-1),\qquad\forall j\in\{21,\ldots,80\},$
$(Z_{21}^{\text{latent}},\ldots,Z_{80}^{\text{latent}})\sim\text{AR}(0.5).$
Conditional on the covariates $Z$, the 20 explanatory variables are generated independent of each other as follows:
$X_{j}\mid(Z_{1},\ldots,Z_{20})\sim\text{Bernoulli}(Z_{20+j}),\qquad\forall j\in\{1,\ldots,20\}.$
Therefore, the knockoffs $\tilde{X}$ are also generated independent of each other as follows:
$\tilde{X}_{j}\mid(Z_{1},\ldots,Z_{20})\sim\text{Bernoulli}(Z_{20+j}),\qquad\forall j\in\{1,\ldots,20\}.$
Appendix A5 Details on the UK Biobank Applications
Background on data and pre-processing.
This study utilizes UK Biobank data (applications 27837 and 74923), including unrelated individuals with complete genotype and phenotype information. Observations recommended for exclusion by the UK Biobank due to insufficient data quality (data field 22010) are omitted. We use ancestries and knockoff variables generated by Sesia et al. (2021). Covariates in all cross-validated lasso models include age (data field 21003), squared age, sex (if applicable, data field 22001), and the top five genetic principal components. The outcome of interest in Section 3.3 is Waist-Hip Ratio (waist circumference /// hip circumference), based on data fields 48 (waist circumference) and 49 (hip circumference).
Background on block-level inference.
As discussed in Appendix A2.3.4 and A2.3.4, Sesia et al. (2021) partition the individual SNPs into contiguous groups at seven levels of resolution based on complete linkage hierarchical clustering. The group sizes range from single SNPs to 425-kilobase (kb) wide groups. For the real data analysis we choose the resolution with median group width 81 kb and for the simulation we choose a resolution with median group width 3 kb. In the simulation, for simplicity, we only choose a single variant within each of these groups to be nonzero. Since inference is based on the block level, the presence of a single nonzero SNP within a block renders the conditional independence hypothesis to be false at the block level as well.
Computational details.
After the construction of knockoff variables and the assembly of cloaked data, the entire computational pipeline for the real data application in Section 3.3 can be run within approximately 15 hours. All analyses are conducted in R/4.2.0 and use the “bigstatsr” and “bigsnpr” packages (Privé et al., 2018) for handling genetic data.
Appendix A6 Additional Simulations with the UK Biobank Data
A6.1 Varying the Proportion of Global Effects
All simulations in this section follow the same general setting as described in Section 3.2. In Figure A4, we fix a signal amplitude of 15, which approximately corresponds to a heritability of 0.5, and vary the proportion of global effects. Below we compare the performance of the aLKF with benchmark methods as described in Section 3.2, as well as the separate subgroups that the Fixed-LKF is based on. Additionally, we include an “unweighted” version of the LKF, where the multiplicative factor $c\in(0,1]$ applied to the lasso penalty is equal to $1$ for all coefficients (uw-aLKF).
Figure A4: Performance of the local knockoff filter (aLKF) and benchmark methods on real genotype data with a simulated outcome while varying the proportion of global effects for a fixed signal amplitude of 15. Group 1 indicates the subgroup $Z_{1}=0,Z_{2}=0$, group 2 indicates $Z_{1}=0,Z_{2}=1$, group 3 indicates $Z_{1}=1,Z_{2}=0$, and group 4 indicates $Z_{1}=1,Z_{2}=1$. The nominal false discovery rate (FDR) level is 0.1. 1% of SNPs are nonzero. 100 iterations.
Figure A4 shows that all methods control the FDR. As expected, for lower proportions of global effects, and, in turn, higher proportions of local effects, the aLKF methods are more powerful than the global approach.
For high proportions of main effects, KF-Global has highest power as expected. However, our objective extends beyond merely identifying main effects - we aim to uncover local effects as well. While KF-Global is able to also reject some of the variants that are locally nonzero, it is not able to pinpoint to the relevant subgroup. Moreover, while the informativeness metric (homogeneity) is close to one for the aLKF methods, it is only increasing to one for higher proportions of global effects for KF-Global.
The aLKF is able to recuperate some of the effects lost to the uw-aLKF with higher proportions of main effects. Overall, the aLKF is able to interpolate between the uw-aLKF for lower proportions of global effects and the KF-Global for higher proportions of global effects.
Figure A5 below shows the results of the same simulation setting for signal sparsity levels of 0.5% and 2%. The results are qualitatively similar.
Figure A5: Simulation setting as in Figure A4 with different levels of sparsity. Sparsity as indicated in row headings.
The next two figures below repeat the same simulation setting, however, the simulations use $Z_{1}$ as the only covariate. Again, the results are qualitatively similar as discussed above.
Figure A6: Same general simulation setting as A4, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.
Figure A7: Same simulation setting as  A6 with different levels of sparsity.
A6.2 Varying the Signal Strength
We include additional simulations in the spirit of Figure 2 in Section 3.2: we vary signal amplitude for a fixed proportion of global effects of 50%. Figure A8 is identical to Figure 2, but includes the uw-aLKF and the separate groups that the Fixed-LKF is based on.
Figure A8: Identical to Figure 2, with the difference being the inclusion of the uw-aLKF, groups 1 to 4 and heritability.
Since there are only local effects in the group $Z_{1}=0$, the subgroups $Z_{1}=0,Z_{2}=0$, $Z_{1}=0,Z_{2}=1$ have higher power compared to the groups $Z_{1}=1,Z_{2}=0$, $Z_{1}=1,Z_{2}=1$. The next figure displays results for sparsity levels of 0.5% and 2%. The results are consistent with those discussed earlier.
Figure A9: Same simulation setting as in Figure A8, except that different levels of sparsity are used, as indicated in row headings.
The next two figures repeat the simulations above, but use $Z_{1}$ as the only covariate.
Figure A10: Same simulation setting as in A8, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.
Figure A11: Same simulation setting as in Figure A10, except that different levels of sparsity are used, as indicated in row headings.
Appendix A7 Additional Results from the Analysis of WHR
Figure A12: Histogram of the knockoff importance statistics for rejected hypotheses. Local rejections are indicated in blue and global rejections in gray.
Figure A13: Partition agnostic Venn diagram comparing the rejections of the aLKF, KF-Global and the Fixed-LKF (considering only unique loci). There are 536 unique loci rejected for the aLKF (541 total rejections, 10 are rejected in both males and females, i.e. 536 unique loci). The totals for the aLKF in this Venn diagram add up to 536 ($277+215+29+15=536$). Similarly, there are 314 unique loci rejections for the Fixed-LKF method ($215+15+18+66=314$). KF-Global only makes unique loci rejections, therefore the total based on this Venn diagram is 595 as well ($277+215+18+85=595$).
Fixed-LKF
Female only Male only Both Male and Female Not detected
aLKF Entire population 127 36 24 297 484
Female only 35 0 0 6 41
Male only 0 4 0 2 6
Both Male and Female 2 1 1 1 5
Not detected 68 16 0 (84)
232 57 25 (306) 314;536
Table A3: Comparison of the rejections from the aLKF to those from the Fixed-LKF method for unique loci. For the aLKF, 10 loci are rejected twice. There are are 541 total rejections for the aLKF in total, which means that $541-10/2=536$ are unique. The total number of aLKF rejections in this table adds up to 536 ($484+41+6+5=536$). Similarly, 314 of the Fixed-LKF rejections are unique. The total number of unique rejections (314; 536) do not include the totals for “Not detected” loci. The numbers in this table also match with those in the Venn diagram in Figure A13.
{ThreePartTable}{TableNotes}
Local rejections for WHR on the British unrelated population, sorted by the knockoff importance statistic $W$. FDR level: $\alpha=0.1$.
“SNP lead” is the individual SNP within the rejected group of variants with the largest knockoff importance statistic. “BP lead” is the basepair (BP) of “SNP lead”. The minimum and maximum BP for the group are given by “min BP” and “max BP”. “Partition” indicates the partition of the rejection, i.e. whether the local rejection was made for males or females. The column “rej. global” indicates whether the group was rejected by KF-Global. If “rej. global” is blank, then the group was rejected by KF-Global, if it is “no”, then it was not rejected.
“Gene” indicates the gene(s) that the leading SNP or group overall is closest to. These genes have been found using the NHGRI-EBI Catalog (Sollis et al., 2023), FAVOR (Zhou et al., 2023), and the CVDKP (Costanzo et al., 2023). “Previous discovery” indicates whether each gene has been (locally) associated with WHR or a closely related phenotype, WHR adjusted for BMI, before. The entry “local diff.” means that the gene has been found to have differential effects by sex in previous research (Funkhouser et al., 2020; Heid et al., 2010; Shungin et al., 2015; Randall et al., 2013; Winkler et al., 2015). In contrast, the entry “global” means that previous studies have reported a (significant) association between the gene and WHR, or WHR adjusted for BMI, based on the NHGRI-EBI Catalog and/or the CVDKP (Sollis et al., 2023; Costanzo et al., 2023). If the entry “previous discovery” is missing, then the gene was reported to be associated with WHR (or WHR adjusted for BMI) according to the CVDKP (Costanzo et al., 2023), but not significantly.
Table A4: Validation and annotation of local discoveries
CHR Group W SNP lead BP lead min BP max BP partition rej. vanilla gene previous discovery
$6$ $735$ $0.00496$ rs72961013 $127,529,780$ $126,504,799$ $127,530,850$ female RSPO3 local diff.
$6$ $735$ $0.00121$ rs72961013 $127,529,780$ $126,504,799$ $127,530,850$ male RSPO3 local diff.
$2$ $973$ $0.00314$ rs79953491 $165,555,539$ $165,501,849$ $165,829,017$ female COBLL1 local diff.
$6$ $371$ $0.00305$ rs4711750 $43,757,082$ $43,757,082$ $43,781,321$ female VEGFA local diff.
$3$ $807$ $0.00234$ rs9833879 $129,333,910$ $129,097,692$ $129,637,348$ female PLXND1 local diff.
$3$ $496$ $0.00193$ rs4132228 $64,708,114$ $64,692,129$ $64,862,311$ female ADAMTS9-AS2 local diff.
$1$ $997$ $0.00170$ rs6427245 $170,303,073$ $170,171,258$ $170,345,383$ female GORAB-AS1, LINC01681 local diff.
$9$ $595$ $0.00159$ rs10991417 $107,688,629$ $107,676,940$ $107,743,189$ female ABCA1 local diff.
$5$ $365$ $0.00155$ rs3936510 $55,860,866$ $55,828,080$ $55,880,301$ female MAP3K1, C5orf67 local diff.
$12$ $824$ $0.00151$ rs4765127 $124,460,167$ $124,256,619$ $124,499,584$ female CCDC92, RFLNA, ZNF664 local diff.
$4$ $557$ $0.00146$ rs1377290 $89,719,368$ $89,468,618$ $89,744,242$ female FAM13A local diff.
$1$ $1,006$ $0.00135$ rs9286854 $172,366,806$ $172,302,544$ $172,463,933$ female DNM3 global
$1$ $1,283$ $0.00126$ rs4846567 $219,750,717$ $219,622,746$ $219,979,718$ female ZC3H11B, LYPLAL1-AS1 local diff.
$12$ $258$ $0.00119$ rs718314 $26,453,283$ $26,440,451$ $26,532,354$ female ITPR2 local diff.
$11$ $402$ $0.00114$ rs35169799 $64,031,241$ $63,858,714$ $64,219,679$ female PLCB3 global
$3$ $164$ $0.00110$ rs4684859 $12,498,401$ $12,477,055$ $12,859,500$ female PPARG,TSEN2, CAND2 local diff.
$3$ $164$ $0.00027$ rs2305398 $12,856,856$ $12,477,055$ $12,859,500$ male PPARG,TSEN2, CAND2 local diff.
$11$ $393$ $0.00107$ rs72927213 $62,351,198$ $62,309,586$ $62,649,463$ female TUT1 global
$6$ $598$ $0.00106$ rs2503097 $100,606,568$ $100,599,675$ $100,637,297$ female MCHR2-AS1,SIM1,LOC728012 local diff.
$5$ $678$ $0.00103$ rs17145265 $118,698,998$ $118,676,043$ $118,969,960$ female no TNFAIP8 local diff.
$7$ $244$ $0.00100$ rs1055144 $25,871,109$ $25,852,328$ $25,938,593$ female LINC03007,MIR148A local diff.
$2$ $945$ $0.00099$ rs55920843 $158,412,701$ $158,390,468$ $158,487,634$ female ACVR1C global
$4$ $245$ $0.00090$ rs2048508 $26,079,994$ $25,990,084$ $26,123,616$ female LINC02357, RBPJ global
$6$ $298$ $0.00087$ rs114344942 $34,207,990$ $34,060,867$ $34,238,514$ female HMGA1 local diff.
$2$ $707$ $0.00083$ rs17041868 $111,894,720$ $111,810,633$ $112,007,943$ female BCL2L11,MIR4435-2HG global
$3$ $163$ $0.00081$ rs4279078 $12,314,846$ $12,029,991$ $12,475,557$ female SYN2,PPARG local diff.
$12$ $820$ $0.00079$ rs73228032 $123,322,415$ $123,159,176$ $123,387,922$ female no HIP1R, HCAR1 local diff.
$8$ $748$ $0.00076$ rs6982636 $126,479,315$ $126,447,844$ $126,543,098$ female no TRIB1AL global
$20$ $292$ $0.00075$ rs143384 $34,025,756$ $33,886,799$ $34,797,801$ male GDF5 local diff.
$11$ $364$ $0.00074$ rs12361415 $47,474,146$ $47,370,041$ $49,425,463$ male CELF1,RAPSN, AGBL2 global
$11$ $364$ $0.00022$ rs12286721 $47,701,528$ $47,370,041$ $49,425,463$ female CELF1,RAPSN, AGBL2 global
$10$ $815$ $0.00067$ rs2254069 $122,875,589$ $122,816,108$ $123,000,561$ female WDR11,RPL19P16, MIR5694 local diff.
$12$ $825$ $0.00067$ rs12829378 $124,516,911$ $124,504,283$ $124,637,099$ female no RFLNA global
$17$ $249$ $0.00060$ rs16961845 $17,432,456$ $17,432,456$ $17,693,891$ female PEMT global
$14$ $87$ $0.00055$ rs12432376 $25,942,829$ $25,920,720$ $25,990,890$ male no STXBP6 local diff.
$1$ $872$ $0.00053$ rs114056923 $154,868,744$ $154,824,330$ $155,019,710$ female no PMVK,KCNN3 global
$1$ $528$ $0.00049$ rs111573261 $66,001,886$ $65,896,561$ $66,169,679$ female LEPR global
$3$ $932$ $0.00047$ rs62271364 $150,050,596$ $150,004,071$ $150,229,958$ female no TSC22D2,LOC646903, AC018545.1 local diff.
$10$ $608$ $0.00046$ rs1236816 $89,684,214$ $89,591,344$ $89,850,500$ female PTEN global
$1$ $252$ $0.00046$ rs649458 $23,712,528$ $23,294,671$ $23,829,648$ female TCEA3 global
$1$ $873$ $0.00042$ rs74593044 $156,016,868$ $155,025,361$ $156,022,802$ female no UBQLN4 global
$1$ $799$ $0.00041$ rs72694053 $115,391,936$ $114,931,419$ $115,672,145$ male SYCP1 global
$5$ $335$ $0.00037$ rs72765634 $50,799,239$ $50,420,111$ $50,950,593$ male no HMGB1P47,RNA5SP182 global
$3$ $853$ $0.00034$ rs2306374 $138,119,952$ $137,784,095$ $138,565,389$ female no MRAS, ESYT3 local diff.
$3$ $853$ $0.00017$ rs808925 $138,177,119$ $137,784,095$ $138,565,389$ male no MRAS, ESYT3 local diff.
$2$ $755$ $0.00030$ rs954244 $121,309,231$ $121,305,604$ $121,355,408$ female no LINC01101, Y_RNA global
$2$ $421$ $0.00029$ rs4953618 $48,954,842$ $48,954,842$ $49,011,274$ female no GTF2A1L,LHCGR,STON1-GTF2A1L global
$2$ $375$ $0.00026$ rs116520905 $44,037,396$ $43,937,198$ $44,049,795$ female no ABCG5, DYNC2LI1 global
$1$ $19$ $0.00026$ rs12044475 $2,709,392$ $2,373,987$ $2,802,562$ male no TTC34
$1$ $396$ $0.00025$ rs77213763 $46,440,828$ $45,782,786$ $46,897,679$ male MAST2,AL358075.4,PIK3R3
$1$ $396$ $0.00020$ rs785493 $46,584,859$ $45,782,786$ $46,897,679$ female MAST2,AL358075.4,PIK3R3
$2$ $251$ $0.00024$ rs73920682 $25,578,778$ $25,487,656$ $25,592,436$ female no DNMT3A,AC012074.1 global
$2$ $135$ $0.00021$ rs4799 $10,933,249$ $10,881,867$ $10,969,662$ female no PDIA6 global
$1$ $105$ $0.00020$ rs2071931 $9,329,289$ $9,305,445$ $9,339,467$ female H6PD local diff.
$1$ $99$ $0.00018$ rs301805 $8,481,016$ $8,374,253$ $8,982,906$ male no RERE global
$7$ $283$ $0.00017$ rs12112389 $30,936,024$ $30,932,526$ $30,942,744$ female no MINDY4,AQP1 global
$16$ $669$ $0.00017$ rs9926861 $85,156,713$ $85,150,638$ $85,172,619$ female no LINC02139,FAM92B global
Appendix A8 The Robust Local Knockoff Filter
A8.1 Testing Partial Conjunction Local Hypotheses
Recall that the local knockoff filter is designed to test hypotheses (5) of the type
$\mathcal{H}_{j,\ell}(\nu)=\bigcap_{i\in[n]\,:\,\nu_{j}(Z^{i})=\ell}\mathcal{H}^{i}_{j},\qquad j\in[p],\ell\in[L_{j}],$
for a (possibly random) partition function $\nu$ with total width $L=\sum_{j=1}^{p}L_{j}$. Here, we present an extension of our method, inspired by Li et al. (2022), that makes it possible to test the following family of partial conjunction (Friston et al., 2005; Benjamini and Heller, 2008) hypotheses:
$\mathcal{H}^{\mathrm{pc},r}_{j}(\nu):\left|\left\{\ell\in[L_{j}]:\mathcal{H}_{j,\ell}(\nu)\text{ is true}\right\}\right|>L_{j}-r,$ (A12)
for any fixed $r\in\{1,\ldots,L_{j}\}$ and any $j\in[p]$. Intuitively, the hypothesis in (LABEL:eq:null-pici) states that the number of subgroups for variable $j$ in which $\mathcal{H}_{j,\ell}(\nu)$ is false is strictly smaller than $r$. Therefore, rejecting (LABEL:eq:null-pici) suggests $X_{j}$ is associated with $Y$ in at least $r$ distinct subgroups identified by the partition function $\nu$.
If $r=1$, the partial conjunction hypothesis becomes equivalent to a global hypothesis; in that case, rejecting $\mathcal{H}^{\mathrm{pc},r}_{j}(\nu)$ suggests the existence of least one subgroup in which $X_{j}$ is conditionally associated with $Y$. By contrast, rejecting $\mathcal{H}^{\mathrm{pc},r}_{j}(\nu)$ for $r=G$ requires evidence of a conditional association between $X_{j}$ and $Y$ across all subgroups identified by the partition function $\nu$.
[…]
For each $j\in\{1,\ldots,p\}$, let $n^{-}_{j}$ count the negative signs among the $W_{j,\ell}$ statistics for all subgroups $\ell\in[L_{j}]$, and let $\smash{n^{0}_{j}}$ be the number of zeros among them. For simplicity, assume $\hat{\nu}$ partitions the covariate space into $L_{0}$ subgroups for each variable; this is without loss of generality, as it is otherwise sufficient to set the remaining undefined statistics to zero if $L_{j}<L_{0}$ for some $j\in[p]$. Then, compute
$\begin{split}p_{j}^{\mathrm{pc},r}&=\Psi\left(n^{-}_{j}-1,(L_{0}-r+1-n^{0}_{j})\lor 0,\frac{1}{2}\right)+U_{j}\cdot\Psi^{\prime}\left(n^{-}_{j},(L_{0}-r+1-n^{0}_{j})\lor 0,\frac{1}{2}\right),\end{split}$ (A13)
where $\Psi(\cdot,m,\pi)$ is the binomial cumulative distribution function, $\Psi^{\prime}(\cdot,m,\pi)$ is the corresponding probability mass, and $U_{j}$ is a uniform random variable on $[0,1]$ independent of everything else. Further, define
$|W_{j}^{\mathrm{pc},r}|=\bar{w}\left(|W_{j,1}|,\dots,|W_{j,L_{0}}|\right),$ (A14)
for some symmetric function $\bar{w}$, such as that which multiplies the top $r$ largest entries by absolute value:
$w\left(|W_{j,1}|,\dots,|W_{j,L_{0}}|\right)=\prod_{\ell=1}^{r}\bar{|W|}_{j,(L_{0}-\ell+1)}.$
Above, $\bar{|W|}_{j,(\ell)}$ are the order statistics for $\{|W_{j,1}|,\dots,|W_{j,L_{0}}|\}$. Then, the FDR for (LABEL:eq:null-pici) can be controlled by applying the usual selective SeqStep+ sequential testing procedure of Barber and Candès (2015), thanks to a relatively simple extension of Theorem 1 from Li et al. (2022).
Theorem A1.
[…]
Proof.
This result follows immediately from Theorem 1 in Li et al. (2022) with a simple extension of Proposition 5 therein. The original statement of Proposition 5 in Li et al. (2022) assumed the statistics $\mathbf{W}$ to be computed using data collected from separate and a-priori fixed experimental settings (or environments), but it is easy to see that their proof only requires the flip-sign property established by our Theorem 2. ∎
Alternatively, the FDR for (LABEL:eq:null-pici) could be (approximately) controlled by applying to the p-values (LABEL:eq:pici-pvalues) ordered by (LABEL:eq:pici-order) the accumulation test (Li and Barber, 2017) instead of SeqStep+; this can be easily proved with an analogous extension of Theorem 2 from Li et al. (2022).
A8.2 Experiments in a “Transfer Learning” Setting
We demonstrate the empirical performance of the Robust-aLKF applied in a “transfer learning” setting inspired by Li et al. (2022). The goal is complementary to that of the experiments in Section 3: we wish to detect variables with a robust association across different subgroups. This is useful for example to find which variables maintain their association with $Y$ within a future data set with covariate shift, in which the distribution of $Z$ differs from the current one but the true model for $Y\mid X,Z$ is the same (Li et al., 2022).
To simulate this covariate shift scenario, we generate data from a model similar to that in Section 3.1, but now only half of the non-null variables interact with the covariates; i.e., $\smash{\beta_{j}^{(i)}}=\bar{\beta}$ for half of the variables. Further, the data dimensions are increased to ensure sufficiently many discoveries can be made. Specifically, the number of variables is $p=40$, while the number of covariates is $m=160$, of which 120 are continuous. The covariate shift is imagined to be such that the 40 binary covariates are always equal to zero instead of following a symmetric Bernoulli distribution. Therefore, the goal is to identify the $20$ variables whose association does not involve interactions, as the covariate shift would make the others irrelevant. These data are analyzed as in Section 3.1, with the only difference that the aLKF is replaced by its robust version outlined in Appendix LABEL:app:methods-ralkf, which is designed to test (LABEL:eq:null-pici).
Figure LABEL:fig:experiment-transfer-1 reports on the results of 100 independent experiments, as a function of the sample size. The performance of the robust version of the aLKF is quantified in terms of the average FDP, the power, and the homogeneity of the reported findings, separately within the training population and under covariate shift. In the latter case, only the discoveries of variables with a direct conditional association with the outcome (i.e., not mediated by any interactions) are counted as true, while the others are considered false positives. Therefore, neither the naive benchmark nor the global knockoff filter control this notion of FDR under covariate shift because they tend to report all associated variables, including those that are non-null only thanks to interactions. By contrast, the Robust-aLKF empirically controls the FDR even under covariate shift. The data-splitting benchmark here is a modified version of our Robust-aLKF in which half of the observations are used for partitioning the covariate space, without randomly swapping the knockoffs, and the remaining half are utilized for testing the selected hypotheses. The results demonstrate data-splitting is valid under covariate shift but it is not as powerful as our method, especially if the sample size is moderately large.

### v2

@@
 A key innovation of the LKF is its ability to test not only pre-defined, fixed hypotheses but also data-driven local hypotheses, automatically identified by powerful machine learning models as likely to be informative for the data at hand.
-This approach introduces a new way to efficiently partition the information in a dataset, allowing both learning and testing (Tian and Taylor, 2018; Rasines and Young, 2023; Leiner et al., 2023).
-Thus, our method can be seen as a “data-thinning” approach (Dharamshi et al., 2024), specifically designed for knockoff inference and taking advantage of the unique capacity of knockoffs to partially “mask” data (Spector and Fithian, 2022).
+This method efficiently partitions the information in a dataset, allowing both learning and testing (Tian and Taylor, 2018; Leiner et al., 2023).
+Thus, it can be seen as a “data-thinning” approach (Dharamshi et al., 2024), specifically designed for knockoff inference and taking advantage of the unique capacity of knockoffs to partially “mask” data (Spector and Fithian, 2022).
 After outlining the general framework of the LKF, which is very flexible, we describe specific implementations, including an approach tailored to the analysis of large-scale GWAS data.
@@
 This paper focuses on identifying subgroups of a heterogeneous population within which there are relatively homogeneous conditional associations.
-There are many algorithms and models that can discover subgroups of individuals exhibiting coherent behaviors (Novak et al., 2009), detect interactions (Bien et al., 2013; Wen and Stephens, 2014; Tibshirani and Friedman, 2020), or estimate heterogeneous effects in regression and causal inference (Athey and Imbens, 2016; Moore et al., 2019; Künzel et al., 2019; Hahn et al., 2020; Cui et al., 2023); see also Caron et al.
+There are many algorithms and models that can discover subgroups exhibiting coherent behaviors (Novak et al., 2009; Chernozhukov et al., 2018), detect interactions (Bien et al., 2013; Wen and Stephens, 2014; Tibshirani and Friedman, 2020), or estimate heterogeneous effects in regression and causal inference (Athey and Imbens, 2016); see also Caron et al.
 (2022) for a review.
-While such methods are usually not designed to obtain statistical inferences with finite-sample guarantees, they provide powerful tools that can be leveraged in the “learning” phase of our method.
-This work is also inspired by the literature on explainability in machine learning; e.g., see Christoph (2020) for a review.
-A central challenge in this field is enhancing the interpretability of opaque predictive models by identifying “important” variables.
+While such methods are usually not designed to obtain statistical inferences with finite-sample guarantees, they can be leveraged in the “learning” phase of our method.
+This work is also inspired by the literature on explainability in machine learning (Molnar, 2020), where a central challenge in this field is enhancing the interpretability of opaque predictive models by identifying “important” variables.
 There are many definitions of variable importance, from descriptive measures capturing a model’s input-output relationship to inferential measures offering insights into the underlying population.
@@
 However, one limitation of this approach, which we address in this paper, is that it does not inform about how the importance of $X_{2}$ varies with changes in the value of $Z_{1}$.
-In other words, testing (1) cannot tells us that $X_{2}$ is only locally important among the subgroup with $Z_{1}=1$.
+In other words, testing (1) cannot tell us that $X_{2}$ is only locally important among the subgroup with $Z_{1}=1$.
 As a first step towards capturing heterogeneity, Li et al. (2022) extended the framework of Candès et al. (2018) to test a family of environment-level hypotheses defined as
@@
 The concept of an individual-specific distribution $p^{i}_{Y\mid X,Z}$ is abstract and provides limited opportunities for learning, as each individual essentially forms their own category.
-We use this framework to capture unknown heterogeneity, but in practice, learning will be conducted at the group level by aggregating information from “similar” individuals, as it will soon become clear.
+We use this framework to capture unknown heterogeneity, but in practice, learning will be conducted at the subgroup level by aggregating information from “similar” individuals, as it will soon become clear.
 Given the distribution (3) one may consider testing a family of individual-level conditional hypotheses defined as:
@@
 2.4 Modeling Assumptions and Knockoffs
-In general, Equation (3) can be factored as $P(\mathbf{X},\mathbf{Y},\mathbf{Z})=\prod_{i=1}^{n}p_{Z}(Z^{i})\cdot p_{X\mid Z}(Z^{i}\mid X^{i})\cdot p^{i}_{Y\mid X,Z}(Y^{i}\mid X^{i},Z^{i})$.
+In general, Equation (3) can be factored as $P(\mathbf{X},\mathbf{Y},\mathbf{Z})=\prod_{i=1}^{n}p_{Z}(Z^{i})\cdot p_{X\mid Z}(X^{i}\mid Z^{i})\cdot p^{i}_{Y\mid X,Z}(Y^{i}\mid X^{i},Z^{i})$.
 Following a model-X approach (Candès et al., 2018), we treat $\mathbf{Z}$ as fixed and $P_{X\mid Z}$ as known, while allowing $p^{i}_{Y\mid X,Z}$ to be arbitrary and unknown for all $i\in[n]$.
@@
 Constructing knockoffs that satisfy (7) not only requires knowledge of $P_{X\mid Z}$ but it can also be computationally expensive in general.
-However, the problem is well-studied (Candès et al., 2018; Gimenez et al., 2019; Bates et al., 2020), with practical and effective solutions available for GWAS data (Sesia et al., 2019, 2020, 2021).
+However, the problem is well-studied (Candès et al., 2018; Gimenez et al., 2019; Bates et al., 2020), with practical and effective solutions available for GWAS data (Chu et al., 2024; Sesia et al., 2019, 2020, 2021).
 Therefore, in this paper we assume the availability of suitable knockoffs for the data at hand, using this as a starting point for developing our methodology to test local conditional independence hypotheses.
+In our GWAS analyses specifically, we follow the approach of Sesia et al.
+(2021) to generate valid knockoffs, leveraging hidden Markov models.
 2.5 A Knockoff Filter for Fixed Local Hypotheses
@@
 More precisely, $\mathbf{T}$ (respectively, $\tilde{\mathbf{T}}$) is the concatenation of $p$ sub-vectors $(T_{j,\ell})_{\ell\in[L_{j}]}$ (respectively, $(\tilde{T}_{j,\ell})_{\ell\in[L_{j}]}$) for $j\in[p]$, where each element indexed by $(j,\ell)$ quantifies the importance of $X_{j}$ (respectively, $\tilde{X}_{j}$) in predicting $Y$ within the subgroup defined by $\nu_{j}(z)=\ell$, according to a suitable predictive model for $Y$ given $(X,\tilde{X},Z)$.
-To enable provable FDR control, these scores must be computed in such a way that swapping a variable $X_{j}$ with its knockoff $\tilde{X}_{j}$ within any subgroup $\ell\in[L_{j}]$ would have the sole effect of swapping the corresponding scores and in that subgroup, namely $T_{j,\ell}$ and $\tilde{T}_{j,\ell}$.
+To enable provable FDR control, these scores must be computed in such a way that swapping a variable $X_{j}$ with its knockoff $\tilde{X}_{j}$ within any subgroup $\ell\in[L_{j}]$ would have the sole effect of swapping the corresponding scores in that subgroup, namely $T_{j,\ell}$ and $\tilde{T}_{j,\ell}$.
 We now formalize this principle within a general framework, which not only encompasses all implementations of our method applied in this paper but also accommodates a range of potential future extensions.
@@
 Let $\mathbf{V}\in\{0,1\}^{n\times p}$ be i.i.d.
-Bernoulli random variables, representing a random noise matrix generated by the analyst independent of everything else.
+Bernoulli(0.5) random variables, representing a random noise matrix generated by the analyst independent of everything else.
 Then, let $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}\in\mathbb{R}^{n\times 2p}$ denote the concatenation of $\mathbf{X}$ and $\tilde{\mathbf{X}}$, with the $i$-th observation of $X_{j}$ swapped with its knockoff if and only if $V_{ij}=1$.
@@
 This model can be trained using the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ and a smaller local dataset $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$, including only observations $(X_{j},\tilde{X}_{j},Y,Z)$ for individuals $i\in[n]$ with $\nu_{j}(Z^{i})=\ell$.
-Figure A2 in Appendix A2 provides a schematic visualization of this workflow.
+Figure 1 provides a schematic visualization of this workflow.
 Then, to guarantee (9), it suffices to compute $(T_{j,\ell},\tilde{T}_{j,\ell})$ in such a way that swapping $X_{j}$ and $\tilde{X}_{j}$ in $\mathcal{D}^{(j,\ell)}$ results in $T_{j,\ell}$ being swapped with $\tilde{T}_{j,\ell}$; see Proposition A3 in Appendix A3.
@@
 In other words, for each $j\in[p]$, we can use any random partition function $\smash{\hat{\nu}_{j}:{\mathcal{Z}}\mapsto{\mathbb{N}}}$ whose parametrization may depend on $\tilde{\mathcal{D}}(\mathbf{V})$; i.e., $\smash{\hat{\nu}_{j}(z;\tilde{\mathcal{D}}(\mathbf{V}))\in[\hat{L}_{j}]}$ for all $z\in\mathcal{Z}$ and some $\hat{L}_{j}\in\mathbb{N}_{+}$.
-This two-step framework for adaptively selecting and testing local hypotheses is outlined by Algorithm 2 and visualized schematically in Figure A2 in Appendix A2.
+This two-step framework for adaptively selecting and testing local hypotheses is outlined by Algorithm 2 and visualized schematically in Figure 1 below.
 We call this approach Adaptive Local Knockoff Filter (aLKF).
+Figure 1: Schematic of the Adaptive Local Knockoff Filter.
+The relevant variables and associated statistics for different local environments are shown in different colors.
+The original variables are shown in darker shades and the knockoffs in lighter shades.
 Input: Knockoff-augmented data set $\mathcal{D}=([\mathbf{X},\tilde{\mathbf{X}}],\mathbf{Y},\mathbf{Z})$; FDR level $\alpha\in(0,1)$;
@@
 Otherwise, $\smash{\hat{\nu}_{j}(z)}=1$ for all $z$ values.
-In general, when dealing with $m$ covariates, this approach is applied using a hyper-parameter $G_{\max}\in\mathbb{N}$ controlling the maximum number of interactions that may be selected per variable (e.g., $G_{\max}=2$).
-As detailed in Appendix A2.2, this implementation can be extended to handle multiple binary covariates as well as non-binary covariates.
+In general, when dealing with $m$ covariates, this approach is applied using a hyper-parameter $G_{\max}\in\mathbb{N}$ controlling the maximum number of interactions that may be selected per variable.
+The choice of $G_{\max}$ does not affect the theoretical validity of our method, but it mediates a trade-off: larger values may reveal more specific (and thus potentially more homogeneous) subgroups, but can also reduce power to reject the associated local hypotheses due to the smaller effective sample size within each subgroup.
+In this paper, we fix $G_{\max}=1$ or $G_{\max}=2$, which achieve a reasonable balance and yield informative results.
+Optimally tuning $G_{\max}$ for a specific dataset is an interesting direction for future research.
+As detailed in Appendix A2.2, this implementation can be extended to handle multiple binary covariates as well as non-binary ones.
+For non-binary covariates, we apply discretization, as the ability to detect meaningful subgroups and reject the corresponding null hypotheses generally depends on having a sufficient number of individuals in each subgroup.
 Additionally, Appendix A2.3 describes a specialized implementation with computational shortcuts designed for GWAS-like settings, characterized by a very large number $p$ of variables and a modest number $m$ of covariates.
 3 Empirical Demonstrations
-We demonstrate the performance of aLKF in two simulated settings and one data analysis.
+We demonstrate the performance of aLKF through two simulated settings and one real data analysis.
+The simulation settings are designed to capture the most practically relevant scenarios.
+The first, presented in Section 3.1, investigates variable selection across many environments ($m$) with a moderate number of variables ($p$).
+The second, described in Section 3.2, reflects GWAS-like conditions with high dimensionality ($p>n$) and fewer environments, serving as a stepping stone to the real GWAS data analysis in Section 3.3.
 3.1 Simulation: Finding Highly Localized Important Variables
 We generate synthetic data with $p=20$ binary variables and $m=80$ covariates, where the first 20 covariates are binary and the remaining 60 are continuous in $[0,1]$.
-Details on the joint distribution of these variables and the construction of knockoffs are in Appendix A4.
+Details on the joint distribution of these variables are in Appendix A4.
 The outcome is generated using a “causal” linear model with heterogeneous effects and homoscedastic standard Gaussian noise.
@@
 Among the remaining 40 covariates, half (randomly chosen) have effects with absolute value 4 and independent random signs, while the others have no effects; i.e., $\gamma_{41},\ldots,\gamma_{80}\in\{0,\pm 4\}$.
+Based on this simulation setup, valid knockoffs can be sampled independently conditional on the covariates from a Bernoulli distribution; see Appendix A4 for further details.
 Given a sample of $n$ independent observations, our goal is to identify which variables are non-null within which subgroups, aiming for both power (many rejections) and precision (informative subgroups).
 We apply our method (aLKF) at the nominal FDR level $\alpha=0.1$, using the lasso-based approach for importance scores (Appendix A2.1) and learning data-driven partitions via a lasso model with interactions (Appendix A2.2), allowing at most two interactions per variable ($G_{\max}=2$).
-We compare aLKF against three benchmarks: (1) Global-KF, the knockoff filter of Candès et al.
+We compare aLKF against three benchmarks that are applicable within the knockoffs framework: (1) Global-KF, the knockoff filter of Candès et al.
 (2018), which controls the FDR but tests only population-wide hypotheses; (2) Split-LKF, a less powerful version of aLKF that splits the data, using half to learn the partition function and the other half to test local hypotheses, sacrificing power due to reduced sample size but maintaining FDR control; (3) Naive-LKF, an invalid version of aLKF that uses all data twice—to learn the partition and to test local hypotheses—without cloaking, failing to control the FDR due to selection bias.
@@
 3.2 Simulation: Searching for Gene-Environment Interactions
-The following experiments simulate a search for gene-environment interactions in GWAS data, using the real chromosome-21 genotypes of unrelated white non-British subjects in the UK Biobank.
-We use the knockoffs defined by Sesia et al.
-(2021).
+The following experiments simulate a search for gene-environment interactions in GWAS data, using real chromosome 21 genotypes from unrelated white non-British individuals in the UK Biobank.
+Working with real genetic variants introduces additional complexity due to strong collinearity among nearby variants on the same chromosome—a phenomenon known as linkage disequilibrium (LD).
+Because of LD, identifying the precise causal variant is often infeasible; instead, associations are typically detected at the level of broader genomic regions comprising multiple correlated SNPs.
+Following Sesia et al.
+(2021), we partition the genome into contiguous groups of SNPs in strong LD, which serve as the units of inference in place of individual variants.
+Intuitively, our method tests whether each group of SNPs is conditionally independent of the outcome within a given environment, conditional on all other SNPs and covariates.
+The formal definition of the local hypotheses for groups of SNPs is provided in Appendix A2.3.4.
+The choice of SNP group size reflects a well-known trade-off between resolution and power: smaller groups enable more precise localization of associations but tend to reduce power due to strong correlations among nearby variants (Sesia et al., 2020, 2021; Gablenz and Sabatti, 2025).
+Sesia et al.
+(2021) introduce a hierarchy of LD-based groupings to support multi-resolution analyses, and Gablenz and Sabatti (2025) propose a method to aggregate results across resolutions while controlling the FDR.
+To maintain focus and clarity, we adopt a single resolution in this paper, with a median SNP group width of 3 kilobases (kb)—a scale that captures strong linkage disequilibrium while allowing variation in signal strength.
+Importantly, the group size does not affect the theoretical validity of our method.
+Further details on SNP grouping are provided in Appendices A2.3.4 and A5.
+Because our hypotheses target groups of SNPs, we can use more powerful knockoffs that satisfy a weaker version of the exchangeability condition in Equation (7), which now only needs to hold under group-wise swapping.
+Using knockoffs designed for individual SNPs would reduce power unnecessarily.
+This issue has been addressed by Sesia et al.
+(2021), who developed an algorithm for generating valid knockoffs for SNP groups in GWAS data using hidden Markov models.
+We adopt their knockoffs in our analysis.
 There are $14,733$ individuals in total and $p=8,832$ variants.
@@
 For global effects, this adjustment factor is 1, while for local effects it is approximately $\sqrt{2}$.
-To maintain sufficient power in the presence of strong collinearity among nearby genetic variants on the same chromosome (linkage disequilibrium), we adopt the approach of Sesia et al.
-(2020, 2021).
-The genome is partitioned into contiguous blocks of variants in strong linkage disequilibrium, with conditional hypotheses tested at the block level.
-These blocks serve as the units of inference, replacing individual SNPs.
-In this simulation, the median width of the blocks is 3 kilobases (kb); see Appendices A2.3.4 and A5 for further details.
-Our goal is to identify which variables are non-null within which subgroups, seeking both high power and precision.
-We apply aLKF as described in Section 3.1, allowing the data-driven partition function to select at most one interaction per variable ($G_{\max}=1$).
-In addition to the three benchmarks from Section 3.1Global-KF, Split-LKF, and Naive-LKF—we introduce a fourth benchmark, Fixed-LKF, a non-adaptive version of LKF.
+Our goal is to identify which groups of SNPs are non-null within which subgroups of individuals, aiming for both high power and precise localization.
+We apply aLKF as described in Section 3.1, with modifications to account for SNP grouping as detailed in Appendix A2.3.4, and we constrain the data-driven partition function to select at most one interaction per variable ($G_{\max}=1$).
+In addition to the three benchmarks from Section 3.1Global-KF, Split-LKF, and Naive-LKF—we introduce a fourth, Fixed-LKF, a non-adaptive version of LKF described below.
+All methods in this comparison use the same SNP group definitions and the same group knockoffs from Sesia et al.
+(2021).
 Fixed-LKF applies Algorithm 1 using a fixed partition function with four local environments for each variable: $(Z_{1},Z_{2})\in\{0,1\}^{2}$.
@@
 3.3 Data Analysis: Genetic Underpinning of WHR Across Sexes
-We study the genetic basis of Waist-Hip Ratio (WHR) across sexes using the UK Biobank data (Bycroft et al., 2018); see Appendix A5 for details.
+We study the genetic basis of Waist-Hip Ratio (WHR) across sexes using the UK Biobank data (Bycroft et al., 2018).
 Previous studies have examined sex-specific effects for WHR or WHR adjusted for BMI, identifying between 7 and 44 sex-specific loci (Heid et al., 2010; Randall et al., 2013; Winkler et al., 2015; Shungin et al., 2015; Funkhouser et al., 2020).
@@
 We define WHR = waist circumference / hip circumference.
+After the construction of knockoff variables with the approach of Sesia et al.
+(2021) and the assembly of cloaked data, the analysis described below takes approximately 15 hours on a computing cluster.
+We refer to Appendix A5 for additional computational details.
 We apply aLKF using the specialized implementation described in Appendix A2.3, designed for large-scale GWAS data with a very large number $p$ of explanatory variables.
@@
 The target FDR level is $\alpha=0.1$.
-Similar to the simulation in Section 3.2 and as discussed in Appendix A2.3.4 and A5, inference is conducted using blocks of SNPs, with median width equal to 81 kilobases (kb), as the units of inference.
-To learn an informative partition function, after pre-screening, we fit a lasso model including all possible pairwise interactions between each of the pre-screened SNPs and sex; see Appendix A2.3.4 for further details on the process.
-This leads to the definition of 15,001 block-level hypotheses for the full sample and 175 hypotheses each for males and females.
+As in the simulation in Section 3.2, and detailed in Appendices A2.3.4 and A5, inference is performed at the level of SNP groups, with a median group width of 81 kilobases (kb).
+In real data, this resolution offers a practical balance between power and precision of the tested hypotheses (Sesia et al., 2020, 2021; Gablenz and Sabatti, 2025).
+To learn an informative partition function, after pre-screening, we fit a lasso model including all possible pairwise interactions between the pre-screened SNPs and sex; see Appendix A2.3.4 for further details.
+This leads to the definition of 15,001 SNP-group-level hypotheses for the full sample and 175 hypotheses each for males and females.
 As shown in Table 1, aLKF yields 541 rejections, of which 57 are sex-specific (46 for females and 11 for males).
@@
 The data analysis and simulations presented in this paper all reflect the results of one knockoff generation, with associated randomness.
-In practical applications, one might want to resort to multiple knockoff to increase the stability of selection (Ren et al., 2023).
+In practice, one may wish to use multiple knockoff realizations to improve stability (He et al., 2021).
+An interesting direction for future research would be to investigate how the de-randomization approach of Ren et al.
+(2023), which leverages e-values (Wang and Ramdas, 2022), could be adapted to our setting.
 The first step of our procedure consists in identifying the relevant “neighborhoods” in which to test for association between $Y$ and each of the $X_{j}$.
 In the examples considered in this paper, we found that the lasso with interactions can be useful in defining a partition $\nu_{j}(Z)$.
-We also noted that different penalization levels for main effects and interactions might be appropriate to navigate the trade-off between the precision of the hypotheses we decided to test with the power to reject them, linked to the size of the subgroup defined by $\nu_{j}(Z)$ (see Appendix A2.3.1).
-A better balance between these might be achieved with a resolution-adaptive procedure as the one described in Gablenz and Sabatti (2024).
+We also noted that different penalization levels for main effects and interactions might be appropriate to navigate the trade-off between the precision of the tested hypotheses with the power to reject them, linked to the size of the subgroup defined by $\nu_{j}(Z)$ (see Appendix A2.3.1).
+A better balance between these might be achieved with a resolution-adaptive procedure as the one described in Gablenz and Sabatti (2025), leveraging e-values.
 The local conditional hypotheses introduced in this paper can be thought as one non parametric definition of “interaction” between explanatory variable $X_{j}$ and covariates $Z$ in influencing outcome $Y$.
@@
 Additionally, it serves to illustrate the flexibility of our local knockoff filter method.
-Recall the high-level ideas of the procedure described in Section 2.5, also visualized schematically in Figure A2.
+Recall the high-level ideas of the procedure described in Section 2.5, also visualized schematically in Figure 1.
 For each pair $(j,\ell)$, we extract the scores $(T_{j,\ell},\tilde{T}_{j,\ell})$ from a local predictive model $\hat{\psi}^{(j,\ell)}$ trained using information from the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ and the local dataset $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$, which includes only observations $(X_{j},\tilde{X}_{j},Y,Z)$ for individuals $i\in[n]$ with $\nu_{j}(Z^{i})=\ell$.
 We now outline this implementation in more detail.
-Figure A2: Schematic of the Adaptive Local Knockoff Filter.
-The relevant variables and associated statistics for different local environments are shown in different colors.
-The original variables are shown in darker shades and the knockoffs in lighter shades.
 First, a sparse generalized linear model (e.g., the lasso) $\hat{\psi}^{(0)}$ is trained to predict $\mathbf{Y}$ given the data in $\tilde{\mathcal{D}}(\mathbf{V})$, tuning the regularization via cross-validation.
@@
 Figure A2: Schematic of Adaptive Local Knockoff Filter, focusing on the computation of the statistics for the first (blue) and fifth (green) local environments.
-The test statistics are based on a simple empirical prior for the two variables that is learnt from the cloaked data set with randomly swapped knockoffs.
-Other details are as in Figure A2.
+The test statistics are based on a simple empirical prior for the two variables that is learned from the cloaked data set with randomly swapped knockoffs.
+Other details are as in Figure 1.
 A2.2 Learning Local Environments Using Interaction Models
@@
 Algorithms 4 and 5 summarize this specialized implementation of our method, with additional details provided in Appendices A2.3.1A2.3.3.
-Furthermore, Appendix A2.3.4 describes how our method can be applied after partitioning the genome into blocks of SNPs.
+Furthermore, Appendix A2.3.4 describes how our method can be applied after partitioning the genome into groups of SNPs.
 This accounts for strong linkage disequilibrium (LD) among nearby variants on the same chromosome, enabling a multi-resolution analysis (Sesia et al., 2020, 2021).
@@
 (2018) at different levels of resolution.
-Specifically, they partitioned the genome into contiguous blocks of variants in strong linkage disequilibrium and tested for conditional independence with the phenotype at the block level, using these blocks as the units of inference rather than individual SNPs.
-The same strategy can be seamlessly integrated with our aLKF method.
-Starting with a pre-defined partition $\mathcal{B}$ of the genome into contiguous blocks of SNPs $b\in\mathcal{B}$ at the desired resolution, along with corresponding knockoff genotypes, as described in Sesia et al.
+Specifically, they partitioned the genome into contiguous groups of variants in strong linkage disequilibrium and tested for conditional independence with the phenotype at the level of groups of SNPs.
+These groups of SNPs are then used as the units of inference rather than individual SNPs.
+Specifically, let $\mathcal{B}$ be a partition of the genome into contiguous groups of SNPs $b\in\mathcal{B}$. The equivalent environment-level hypothesis at the group level, see Equation (6), is
+¯$\bar{\mathcal{H}}_{b,\ell}(\nu):Y^{(b,\ell)}\perp\!\!\!\perp X^{(b,\ell)}_{b}\mid X^{(b,\ell)}_{-b},Z^{(b,\ell)}\;\;\;\;b\in\mathcal{B},\ell\in[L_{b}].$
+This hypothesis is analogous to the one in Equation (6), except that inference is at the level of groups of SNPS, rather than single SNPs.
+Similarly, the exchangeability is required to hold for groups of SNPs, see also Equation (7):
+[…]
+Sesia et al.
+(2021) construct valid knockoffs for groups of SNPs using hidden Markov models, see Sesia et al.
+(2021) for further details.
+They construct groups of SNPs using complete-linkage hierarchical clustering.
+We can seamlessly integrate the genetic group level strategy with our aLKF method.
+To test the hypothesis described above, starting with a pre-defined partition $\mathcal{B}$ of the genome into contiguous groups of SNPs $b\in\mathcal{B}$ at the desired resolution, along with corresponding knockoff genotypes, as described in Sesia et al.
 (2020) and Sesia et al.
@@
   1. 1.
-    The data cloaking step using the random matrix $\mathbf{V}$ is performed at the block level, assigning the same values of $V_{ij}$ to all SNPs $j$ within a block $b\in\mathcal{B}$.
-This reflects the idea that the units of inference are now blocks of SNPs and not single SNPs.
+    The data cloaking step using the random matrix $\mathbf{V}$ is performed at the level of groups of SNPs, assigning the same values of $V_{ij}$ to all SNPs $j$ within a group of SNPs $b\in\mathcal{B}$.
+This reflects the idea that the units of inference are now groups of SNPs and not single SNPs.
   2. 2.
-    Each block of SNPs $b$ is assigned a single partition function $\hat{\nu}_{b}$, rather than having distinct partition functions for individual SNPs.
+    Each group of SNPs $b$ is assigned a single partition function $\hat{\nu}_{b}$, rather than having distinct partition functions for individual SNPs.
   3. 3.
-    While pre-screening nominally selects a subset of individual variants, since inference is then performed at the block level, the batch-local predictive models are fitted using not only the pre-screened variants but also all other variants in the same block.
+    While pre-screening nominally selects a subset of individual variants, since inference is then performed at the level of groups of SNPs, the batch-local predictive models are fitted using not only the pre-screened variants but also all other variants in the same genetic group.
   4. 4.
-    For each combination of $b$ and $\ell\in\hat{L}_{b}$, a single pair of local scores $(T_{b,\ell},\tilde{T}_{b,\ell})$ is computed by summing the single-SNP scores corresponding to the individual SNPs within that block.
+    For each combination of $b$ and $\ell\in\hat{L}_{b}$, a single pair of local scores $(T_{b,\ell},\tilde{T}_{b,\ell})$ is computed by summing the single-SNP scores corresponding to the individual SNPs within that group.
 Then, the corresponding test statistic is $W_{b,\ell}=T_{b,\ell}-\tilde{T}_{b,\ell}$.
-Finally, the discoveries obtained with our method will be interpreted as highlighting blocks of SNPs displaying significant associations with the outcome within certain subgroups determined by specific covariate values.
+Finally, the discoveries obtained with our method will be interpreted as highlighting groups of SNPs displaying significant associations with the outcome within certain subgroups determined by specific covariate values.
 Appendix A3 Mathematical Proofs
@@
 The outcome of interest in Section 3.3 is Waist-Hip Ratio (waist circumference /// hip circumference), based on data fields 48 (waist circumference) and 49 (hip circumference).
-Background on block-level inference.
+Background on inference at the level of groups of SNPs.
 As discussed in Appendix A2.3.4 and A2.3.4, Sesia et al.
@@
 In the simulation, for simplicity, we only choose a single variant within each of these groups to be nonzero.
-Since inference is based on the block level, the presence of a single nonzero SNP within a block renders the conditional independence hypothesis to be false at the block level as well.
+Note that the presence of a single nonzero SNP within a group renders the conditional independence hypothesis relative to that group of SNPs to be false as well.
 Computational details.
-After the construction of knockoff variables and the assembly of cloaked data, the entire computational pipeline for the real data application in Section 3.3 can be run within approximately 15 hours.
-All analyses are conducted in R/4.2.0 and use the “bigstatsr” and “bigsnpr” packages (Privé et al., 2018) for handling genetic data.
+The GWAS data analyses in Sections 3.2 and 3.3 are conducted in R/4.2.0 and use the “bigstatsr” and “bigsnpr” packages (Privé et al., 2018) for handling large-scale genetic data.
+For data manipulation, transformation and visualization we use the “tidyverse” package (Wickham et al., 2019).
+Computational details on the construction of knockoff variables based on Hidden Markov Models (HMM) can be found in Sesia et al.
+(2021).
+Their software to construct these knockoffs is available at https://msesia.github.io/knockoffgwas/.
+For reference, the construction of the HMM knockoffs in Sesia et al.
+(2021) took $\approx$ 4 d using 10 cores and 80 GB of memory to generate knockoffs on chromosome 1 for the UK Biobank data ($\approx$ 1 million haplotype sequences, 600,000 SNPs, and 600,000 IBD (identical-by-descent segments)).
+In this paper, we run all jobs on a computing cluster and we parallelize when possible.
+For the simulation in Section 3.2, we submit separate jobs for each amplitude, sparsity, proportion of main effects.
+Runtime for 100 iterations varies by the simulation parameters and ranges from around 2 hours to 12 hours, with a median runtime of around 5 hours.
+We run all experiments on the computing cluster with 1 core and 25GB RAM.
+For the real data analysis in section 3.3, we use 20 CPU cores with a max of 200 GB.
+We describe the additional computational steps for GWAS data in detail in Section A2.
+The pre-screening step takes around 2 hours, creating the interactions and running the lasso takes around 5 hours, and the lasso on the screened data takes around 3 hours.
+We submit different jobs to run the local models within each batch, which take around 5 hours.
+In total, the entire process takes around 15 hours.
+Our batch scripts containing additional details available at https://github.com/msesia/i-modelx.
 Appendix A6 Additional Simulations with the UK Biobank Data
@@
 The next two figures below repeat the same simulation setting, however, the simulations use $Z_{1}$ as the only covariate. Again, the results are qualitatively similar as discussed above.
-Figure A6: Same general simulation setting as A4, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.
-Figure A7: Same simulation setting as  A6 with different levels of sparsity.
+Figure A5: Same general simulation setting as A3, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.
+Figure A6: Same simulation setting as  A5 with different levels of sparsity.
 A6.2 Varying the Signal Strength
@@
 The next two figures repeat the simulations above, but use $Z_{1}$ as the only covariate.
-Figure A10: Same simulation setting as in A8, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.
+Figure A9: Same simulation setting as in A7, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.
 Figure A10: Same simulation setting as in Figure A9, except that different levels of sparsity are used, as indicated in row headings.